### v1

# Extended mean field control problems with constraints: The generalized Fritz-John conditions and Lagrangian method

Abstract
This paper studies the extended mean field control problems under dynamic expectation constraints and/or dynamic state-control constraints. We aim to pioneer the establishment of the stochastic maximum principle (SMP) and the derivation of the backward SDE (BSDE) from the perspective of the constrained optimization using the method of Lagrangian multipliers. To this end, we first propose to embed the constrained extended mean-field control (C-(E)MFC) problems into some abstract optimization problems with constraints on Banach spaces, for which we develop the generalized Fritz-John (FJ) optimality conditions. We then prove the stochastic maximum principle (SMP) for C-(E)MFC problems by transforming the FJ type conditions into an equivalent stochastic first-order condition associated with a general type of constrained forward-backward SDEs (FBSDEs). Contrary to the existing literature, we recast the controlled Mckean-Vlasov SDE as an infinite-dimensional equality constraint such that the BSDE induced by the FJ first-order optimality condition can be interpreted as the generalized Lagrange multiplier to cope with the SDE constraint. Finally, we also present the SMP for constrained stochastic control problems and constrained mean field game problems as consequences of our main result for C-(E)MFC problems.
Keywords: Extended mean-field control, dynamic expectation constraints, dynamic state-control constraints, stochastic maximum principle, generalized Fritz-John optimality condition.
1 Introduction
Motivated by wide applications in finance, engineering and management science, stochastic control problems with various types of constraints have attracted tremendous attention over the past decades. Several types of constraints on state process and the control process have been predominantly investigated in different context such as the dynamic (in time) state and/or control constraint, the dynamic (in time) expectation constraint on state and/or control paths, and their simplified versions as static constraints only at the terminal time. New techniques are inevitably needed in order to address the constraints such that the dynamic programming approach and Pontryagin’s maximum principle approach can be modified and adopted.
To name a few in the direction of studies to cope with the state and/or control constraints, Hu and Zhou [27] investigate a stochastic LQ control problem with random coefficients by employing Tanaka’s formula, where the control variable is constrained in a cone; Bonnans and Silva [10] examine the stochastic control problems with control constraints and several equality and inequality constraints over the final state, for which the first order and second order optimality conditions are established; Liu et al. [31] consider stochastic control with an end-point constraint by establishing a general FJ condition in Banach space where only the control variable is reformulated as an optimization problem; Hu et al. [25] investigate a stochastic LQ control problem with regime switching, random coefficients and cone control constraint, where the optimal feedback control and optimal cost value are characterized explicitly via two systems of extended stochastic Riccati equations; Zhang and Zhang [42] study a stochastic linear-quadratic (LQ) control problem with a static equality constraint on the terminal state by using Lagrangian duality method.
On the other hand, fruitful studies to incorporate dynamic or static expectation constraints can also be found in the literature. For example, Soner and Touzi [40] study a stochastic target problem by using the geometric dynamic programming approach; Bouchard and Nutz [13] develop the weak dynamic programming principle for stochastic control problems with expectation constraints that implies the Hamilton-Jacobi-Bellman equation in the viscosity sense; Bokanowski et al. [11] employ the level-set approach to reformulate the constrained control problems into an optimization problem over a family of unconstrained singular control problems; Chow et al. [16] reformulate some state constraints as equality expectation constraints and examine the dynamic programming principle; Pfeiffer et al. [36] develop a duality approach to study the stochastic control problems under both equality and inequality expectation constraints using convex analysis techniques; Hu et al. [26] consider optimal control of stochastic differential equations subject to dynamic expectation constraint where the constrained FBSDEs are formulated and studied; Bayraktar and Yao [3] recently analyze an optimal stopping problem with some inequality and equality expectation constraints in a general non-Markovian framework and establish a dynamic programming principle in weak formulation; Bayraktar and Yao [4] further generalize the methodology in [3] to address stochastic control/stopping problems where the diffusion can be controlled.
MFC problems, also called the McKean-Vlasov optimal control problems, feature the cooperative interactions in the mean field model where all agents jointly optimize the social optimum that yields the optimal control from the perspective of the social planner, which have raised a lot of research interests in the past few years. To name a few, Acciaio et al. [1] explore suitable versions of the Pontryagin stochastic maximum principle (SMP) for extended MFC problems in which the cost function and the state process depend on the joint distribution of the state and the control process; The first-order variation in Bensoussan [5] and the spike-variation method in Peng [35] have been generalized to cope with MFC problems among [2, 7, 33, 21] that handle McKean-Valsoc dynamics with coefficients depending on the moments of the population state distribution; Li [30] examines MFC problems and establish the SMP in scalar interaction forms; Bensoussan et al. [6] introduce a unified approach for MFC and MFG problems by using the HJB-FP coupled equations and the SMP; Carmona and Delarue [14] solve a class of MFC problems with a general dependence on the law of the state process; Pham and Wei [38] investigate the dynamic programming principle for MFC problems in the presence of common noise; Pham and Wei [39] examine the extended MFC with state-control joint law dependence by reformulating the problem into a deterministic control problem with only the marginal distribution of the process as controlled state variable; Bo et al. [8] consider a LQ-MFC problem by proving the Gamma convergence of the optimal controls from the N-player games to the mean-field limit model; Nie and Yan [34] study extended MFC problems under partial information; Bo et al. [9] recently establish the SMP and HJB eqaution for an extended MFC problem with state-control joint law and Poissonian common noise by proposing the extension transformation.
MFC and MFG problems under the dynamic expectation constraints and/or the dynamic state-control constraints are relatively under developed. Grammatico et al. [24] study a deterministic MFC problem in the presence of heterogeneous convex constraints at the level of individual agents; Du and Wu [22] consider a LQG MFC problem with control constraints; Bonnans et al. [12] analyze the existence of equilibria for a class of deterministic MFGs of controls with state and control constraints at the terminal time by using Kahutani’s theorem; Daudin [19] study the convergence problem of mean-field control theory in the presence of state constraints; Daudin [20] examine a problem of optimal control of the Fokker-Planck equation with state constraints in the Wasserstein space and give first-order necessary conditions for optimality in the form of a mean-field game system of partial differential equations associated with an exclusion condition; Craig et al. [18] discuss a class of first-order MFC problem with source and terminal constraints by using a particle method; Meherrem and Hafayed [32] prove Peng’s necessary optimality conditions for a general mean-field system under state constraints based on conventional variational principle.
In the present paper, we study extended mean field control problems under both the dynamic expectation constraints and the dynamic state-control constraints from the perspective of optimization problems with Banach space valued constraints. We propose a new methodology that bridges the SMP method for the C-(E)MFC problems and the Lagrange multipliers method with the generalized Fritz-John (FJ) optimality condition. In a nutshell, when the forward McKean-Vlasov state process is taken as an infinite-dimensional constraints, the generalized FJ optimality condition for the control variable yields the stochastic (first-order) minimum condition that leads to the SMP. Meanwhile, the generalized Lagrangian multiplier in the space of Itô processes induced by the FJ optimality condition on the state variable gives us the adjoint process as the solution to the BSDE.
More concretely, we first develop the type of FJ optimality condition (Theorem 2.7) for nonlinear optimization problems with (Banach space valued) equality and/or inequality constraints by using the generalized gradient introduced in Clarke [17]. Next, as the main novelty of this paper, we formulate our C-(E)MFC problem and transform it into an abstract optimization problem with constraints on Banach spaces (see problem (26)), for which we can apply the newly developed FJ optimality condition on Banach spaces. We then contribute to the establishment of SMP for the C-(E)MFC problem (Theorem 3.11) by transforming the FJ conditions into a stochastic first-order condition that gives rise to a generalized type of constrained FBSDEs. In sharp contrast to the conventional methods, we broaden the scope by considering both state and control as targeted variables in the formulation of the constrained optimization problems such that the Lagrangian multipliers method can be employed. Furthermore, we propose to treat the controlled Mckean-Vlasov state dynamics as an (integral) equality constraint on an infinite dimensional (Hilbert) space, which is also new to the literature. Then, the adjoint process as a solution to a generalized type of constrained BSDEs induced by the FJ first-order optimality condition can be interpreted as the generalized Lagrangian multiplier to handle the forward SDE equality constraint, for which a new space denoted by $\mathbb{M}^{2}$ is introduced, which consists of all Itô processes to characterize the integration formulation of the Mckean-Vlasov type SDE constraint. As special cases of our C-(E)MFC problems, we also derive and present the SMP results for the stochastic optimal control problems as well as mean field game (MFG) problem with the same types of constraints.
The rest of the paper is organized as follows. Section 2 gives an overview of the generalized gradients on Banach space in [17] and establishes the generalized FJ optimality condition and the constraint qualification (CQ) for a class of constrained optimization problems on Banach spaces. Section 3 reformulates our C-(E)MFC as an abstract constrained optimization problem on Banach spaces and develops the SMP and derives the general type of constrained FBSDEs by using the generalized FJ optimality condition. Section 4 presents the SMP results for stochastic optimal control problems and extended MFG problems with the same types of dynamic expectation constraints and state-control constraints as special cases of the main results from Section 3. Finally, the proofs of some auxiliary results in previous sections are reported in Appendix A.
Notations. We list below some notations that will be used frequently throughout the paper:
[…]
||$\cdot$|| ⋅ | Euclidean norm on $\mathbb{R}^{n}$
$(E^{*},\lVert\cdot\rVert_{E^{*}})$ The dual space of a Banach space $(E,\lVert\cdot\rVert_{E})$
$B(x;r)$ The ball centered at $x\in E$ with radius $r>0$
$\langle\cdot,\cdot\rangle_{E,E^{*}}$ The pairing between Banach space $(E,\lVert\cdot\rVert_{E})$ and its dual
$A^{c}$ The complement of the set $A$
$\mathrm{co}(A)$ The convex hull of the set $A\subset E$
$A^{\rm o}$ ($\overline{A}$) The interior (closure) of the set $A\subset E$
$\partial A$ The boundary of the set $A\subset E$
$a\geq_{K}0$ ($a>_{K}0$) $a\in K$ ($a\in K^{\rm o}$) with $K$ being a closed (linear) cone
$Df$ The Fréchet derivative of $f:E\mapsto\mathbb{R}$
$\nabla_{x}f$ Gradient of $f:\mathbb{R}^{n}\mapsto\mathbb{R}$
$D_{X}f$ The partial Fréchet derivative of $f$ with resepct to $X$
$D_{\alpha}f$ The partial Fréchet derivative of $f$ with resepct to $\alpha$
$\mathcal{L}(E_{1};E_{2})$ Set of continuous linear mappings $f:E_{1}\mapsto E_{2}$
$\mathcal{L}(\xi)$ Law of r.v. $\xi$
$C(E_{1};E_{2})$ Set of continuous mappings $f:E_{1}\mapsto E_{2}$
$L^{p}((A,\mathscr{B}(A),\lambda_{A});E)$ Set of $L^{p}$-integrable $E$-valued mapping defined on
measure space $(A,\mathscr{B}(A),\lambda_{A})$. Write $L^{p}(A;E)$ for short
$L^{p}((A,\mathscr{B}(A),\lambda_{A});E)/\mathord{\sim}$ The quotient set of $L^{p}(A;E)$ w.r.t. equivalence
[…]
Write $L^{p}(A;E)/\mathord{\sim}$ for short
$\mathbb{E}$ Expectation operator under probability measure $\mathbb{P}$
$\mathbb{E}^{\prime}$ Expectation operator under probability measure $\mathbb{P}^{\prime}$
${\cal P}_{p}(E)$ Set of probability measures on $E$ with finite $p$-order moments
$\mathcal{M}^{+}[0,T]$ Set of non-negative Radon measures on $[0,T]$
2 Generalized Fritz-John Optimality Condition
For given $m,k\in\mathbb{N}$, denote $I:=\{1,2,\ldots,m\}$ and $J:=\{1,2,\ldots,k\}$ as two index sets. Let $(\mathtt{X},\|\cdot\|_{\mathtt{X}})$ be a Banach space, $(\mathtt{Y}_{i},\|\cdot\|_{\mathtt{Y}_{i}})_{i\in I}$ be a sequence of separable Banach spaces and $(\mathtt{H}_{j},\langle\cdot,\cdot\rangle_{\mathtt{H}_{j}})_{j\in J}$ be a sequence of Hilbert spaces.
In this section, we first develop the optimality condition in the Fritz-John (FJ) type for the following constrained optimization problem on Banach spaces that, for a closed and convex subset $C\subset\mathtt{X}$,
$\begin{cases}\text{minimize $f(x)$ over $x\in C$;}\\ \text{subject to}~{}g_{i}(x)\leq_{K_{i}}0,\quad\forall i\in I,\\ \qquad\qquad~{}~{}h_{j}(x)=0,\quad\forall j\in J,\end{cases}$ (1)
where $K_{i}\subset\mathtt{Y}_{i}$ are closed and convex cones for $i\in I$. The mappings $f:\mathtt{X}\to\mathbb{R}$, $g_{i}:\mathtt{X}\mapsto\mathtt{Y}_{i}$, $h_{j}:\mathtt{X}\to\mathtt{H}_{j}$ for $i\in I$ and $j\in J$ are assumed to be continuously Fréchet differentiable at any $x\in\mathtt{X}$.
For the goal above, we propose some techniques based on generalized gradients introduced in Clarke [17] in which the constraints are only restricted in Euclidean space in the sense that $\mathtt{Y}=\mathtt{H}=\mathbb{R}$ and $K_{i}=\mathbb{R}_{+}\cup\{0\}$ for all $i\in I$. The general Banach space valued constraints formulated in problem (1) are customized to study the C-(E)MFC problems that will be discussed in Section 3.
2.1 A review of generalized gradients on Banach space
For the completeness, let us give the brief overview of the generalized gradient for a locally Lipschitz mapping $f:\mathtt{X}\mapsto\mathbb{R}$ given in Clarke [17]. For any $v\in\mathtt{X}$, we define the generalized directional derivative $f^{\rm o}(x;v)$ in the direction $v$ by
[…]
It follows from Lemma 1 in [17] that $v\mapsto f^{\rm o}(x;v)$ is convex and $|f^{\rm o}(x;v)|\leq c_{f}\lVert v\rVert_{\mathtt{X}}$ with $c_{f}>0$ being the locally Lipschitz coefficient of $f:\mathtt{X}\mapsto\mathbb{R}$. Thus, the generalized gradient of $f:\mathtt{X}\mapsto\mathbb{R}$ at point $x\in\mathtt{X}$ (denoted by $\partial f(x)$) is defined to be the subdifferential of the convex function $v\mapsto f^{\rm o}(x;v)$ at 00. Namely, for $\xi\in\mathtt{X}^{*}$, we have $\xi\in\partial f(x)$ iff $f^{\rm o}(x;v)\geq\langle v,\xi\rangle_{\mathtt{X},\mathtt{X}^{*}}$ for all $v\in\mathtt{X}$. Therefore, one can conclude that $\partial f(x)$ coincides with the subdifferential of $f$ or $\{Df(x)\}$ when $f$ is convex or has a continuous Fréchet derivative $Df:\mathtt{X}\to\mathtt{X}^{*}$, respectively (c.f. Proposition 3 and Proposition 4 in [17]).
Next, we list some basic properties of the generalized gradient whose proofs can be found in [17]:
Lemma 2.1.
The set $\partial f(x)$ satisfies the following properties:
  • (i)
    $\partial f(x)$ is a nonempty, convex and weak- compact subset of $\mathtt{X}^{*}$. Moreover, it holds that $\lVert\xi\rVert_{\mathtt{X}^{*}}\leq c_{f}$ for all $\xi\in\partial f(x)$.
  • (ii)
    For any $v\in\mathtt{X}$, it holds that $f^{\rm o}(x;v)=\max_{\xi\in\partial f(x)}\langle v,\xi\rangle_{\mathtt{X},\mathtt{X}^{*}}$.
  • (iii)
    Let $B$ be a nonempty, convex, and weak- compact subset of $\mathtt{X}^{*}$. Then, $\partial f(x)\subset B$ iff $f^{\rm o}(x;v)\leq\max_{\xi\in B}\langle v,\zeta\rangle_{\mathtt{X},\mathtt{X}^{*}}$ for all $v\in\mathtt{X}$.
  • (iv)
    If $x\in\mathtt{X}$ is a local minimum of $f:\mathtt{X}\mapsto\mathbb{R}$, then $0\in\partial f(x)$.
  • (v)
    Let $\xi_{i}\in\partial f(x_{i})$ with $i\in\mathbb{N}$ and assume that $x_{i}\to x$ and $\xi_{i}\xrightarrow{w^{*}}\xi$ as $i\to\infty$. Then, $\xi\in\partial f(x)$.
  • (vi)
    $\partial(f+g)(x)\subset\partial f(x)+\partial g(x)$ for all $x\in\mathtt{X}$.
We then can show the next result.
Lemma 2.2.
Let $\mathcal{I}$ be an index set and $(\phi_{i})_{i\in\mathcal{I}}$ be a family of uniformly local Lipschitz (real-valued) functions on $\mathtt{X}$ (i.e., their Lipschitz constants are independent of $i\in\mathcal{I}$). Define the mapping $\phi:\mathtt{X}\mapsto\mathbb{R}$ by $\phi(x)=\sup_{i\in\mathcal{I}}\phi_{i}(x)$ for $x\in\mathtt{X}$, and assume that $M(x):=\{i\in\mathcal{I};~{}\phi_{i}(x)=\phi(x)\}\neq\varnothing$. Then, $x\mapsto\phi(x)$ is locally Lipschitz and it holds that
$\partial\phi(x)\subset\overline{\mathrm{co}\{\partial\phi_{i}(x);~{}i\in M(x)\}}.$
Proof.
The first claim is straightforward. It suffices to show the second assertion. Denote by $B$ the nonempty set on the right side of the second assertion. By the uniformly local Lipschitz continuity of $\phi_{i}:\mathtt{X}\mapsto\mathbb{R}$ with $i\in\mathcal{I}$, we conclude that $B$ is a closed, convex and bounded set, and thus it is convex and weak- compact. Furthermore, by Lemma 2.1-(ii), it holds that, for any $v\in\mathtt{X}$,
$\sup_{\xi\in B}\langle v,\xi\rangle_{\mathtt{X},\mathtt{X}^{*}}=\sup_{i\in M(x)}\phi_{i}^{\rm o}(x;v),$
where $\phi_{i}^{\rm o}(x;v)$ is the generalized directional derivative of $\phi_{i}$ in the direction $v$ given in (2), we can directly show from the definition that $\phi^{\rm o}(x;v)$ is no larger than the right-hand side of the above equality. Then, the desired result follows from Lemma 2.1-(iii). ∎
Remark 2.3.
When the index set $\mathcal{I}$ is finite, the right-hand side of the second assertion in Lemma 2.2 can be replaced by $\mathrm{co}\{\partial\phi_{i}(x);~{}i\in M(x)\}$, where $M(x)=\{i\in\mathcal{I};~{}\phi_{i}(x)=\phi(x)\}$ is nonempty (c.f. Proposition 9 in Clarke [17]).
Next, we introduce the distance between a point $x\in C$ and the set $C$ by $d_{C}(x):=\inf_{c\in C}\lVert x-c\rVert_{\mathtt{X}}$. It is easy to see that $x\mapsto d_{C}(x)$ is 1-Lipschitz, and hence the generalized gradient $\partial d_{C}(x)$ is well-defined. Because of the convexity of $C$, for any $x\in C$, the normal cone $N_{C}(x)$ to the set $C$ at the point $x$ can be defined by
$N_{C}(x):=\{\xi\in\mathtt{X}^{*};~{}\langle x-c,\xi\rangle_{\mathtt{X},\mathtt{X}^{*}}\geq 0,~{}\forall c\in C\}.$ (3)
It then follows from Clarke [17] that a point $x\in\mathtt{X}$ minimizes $f$ over $C$ iff there exists a neighborhood $V(x)$ of $x$ such that $x$ minimizes $f+c_{f}d_{C}$ over $V(x)$. Next, for a cone $K_{i}\subset\mathtt{Y}_{i}$ with $i\in I$, let us define
$K_{i}^{+}:=\left\{\xi\in\mathtt{Y}_{i}^{*};~{}\langle y,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\geq 0,~{}\forall y\in K_{i}\right\},\quad\forall i\in I.$ (4)
We have the following dual relationship between $K_{i}\subset\mathtt{Y}_{i}$ and $K_{i}^{+}\subset\mathtt{Y}_{i}^{*}$.
Lemma 2.4.
Let $K_{i}\subset\mathtt{Y}_{i}$ be a closed and convex cone for $i\in I$. It holds that
$K_{i}=\left\{y\in\mathtt{Y}_{i};~{}\langle y,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\geq 0,~{}\forall\xi\in K_{i}^{+}\right\},\quad\forall i\in I.$ (5)
Proof.
For $i\in I$, denote $L_{i}$ as the set on the right hand side of (5). Then, it follows from the definition of $K_{i}^{+}$ that $K_{i}\subset L_{i}$. We will argue by contradiction to show $L_{i}\subset K_{i}$. In fact, if there exists a point $y\in L_{i}$ and $y\notin K_{i}$, the separation theorem yields that, there exists a $\xi\in\mathtt{Y}_{i}^{*}$ such that
$\langle y,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}<0\leq\langle x,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}},\quad\forall x\in K_{i}.$
This leads to $\xi\in K_{i}^{+}$, and hence $\langle y,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\geq 0$ due to the definition of $L_{i}$, which yields a contradiction. ∎
We also recall two important properties on Banach spaces as below.
Lemma 2.5.
For a Banach space $(\mathtt{X},\|\cdot\|_{\mathtt{X}})$, the following results hold:
  • (i)
    If $\mathtt{X}$ is separable, any bounded set in $\mathtt{X}^{*}$ is weak- sequentially compact;
  • (ii)
    If $\mathtt{X}$ is reflexive, any bounded set in $\mathtt{X}^{*}$ is weak- sequentially compact.
We next characterize the interior $K_{i}^{\rm o}$ of the closed and convex cone $K_{i}$ in the next lemma.
Lemma 2.6.
A point $y\in K_{i}^{\rm o}$ iff $\langle y,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}>0$ for all nonzero $\xi\in K_{i}^{+}$.
Proof.
First of all, we assume that $y\in K_{i}^{\rm o}$ and there exists a $\xi\in K_{i}^{+}$ such that $\langle y,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}=0$. As $y\in K_{i}^{\rm o}$, there is a constant $\delta>0$ such that $B(y;\delta)\subset K_{i}^{\rm o}$. Then, for any $\epsilon\in(0,\delta)$ and any $v\in Y_{i}$, thanks to the definition of $K_{i}^{+}$, we can conclude that $\langle y+\epsilon v,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\geq 0$, and hence $\langle v,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\geq 0$, which implies $\xi=0$.
Next, we argue by contradiction to show that the other side of the claim holds. Suppose that $\langle y,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}>0$ for all nonzero $\xi\in K_{i}^{+}$ and $y\notin K_{i}^{\rm o}$, which implies $y\in\overline{K_{i}^{c}}$ with $K_{i}^{c}$ being the complement of $K_{i}$. Consequently, there exists a sequence $(y_{k})_{k\in\mathbb{N}}\subset K_{i}^{c}$ such that $y_{k}\to y$ as $k\to\infty$. Then, in light of Lemma 2.4, there exists a $\xi_{k}\in K_{i}^{+}$ such that $\langle y_{k},\xi_{k}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}<0$ for all $k\in\mathbb{N}$. We assume without loss of generality that $\|\xi_{k}\|_{\mathtt{Y}_{i}^{*}}=1$. Thus, by Lemma 2.5-(i), there exists a weak- convergent subsequence of $(\xi_{k})_{k\in\mathbb{N}}$ (still denoted by $(\xi_{k})_{k\in\mathbb{N}}$) such that $\xi_{k}\to\xi\in\mathtt{Y}_{i}^{*}$ in the topology of weak- convergence and $\xi\in K_{i}^{+}$ results from the definition of $K_{i}^{+}$ Hence, it holds that
$\left|\langle y_{k},\xi_{k}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}-\langle y,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\right|$ $\leq\left|\langle y_{k}-y,\xi_{k}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\right|+\left|\langle y,\xi_{k}-\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\right|$
$\leq\|y_{k}-y\|_{\mathtt{Y}_{i}}+\left|\langle y,\xi_{k}-\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\right|.$
In fact, the second term above converges to zero as $k\to\infty$ in lieu of the definition of weak- convergence. Thus $\langle y,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}=\lim_{k\to\infty}\langle y_{k},\xi_{k}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\leq 0$, which contradicts $\langle y,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}>0$, and the proof is complete. ∎
2.2 The generalized FJ optimality condition on Banach spaces
In this subsection, we establish a necessary optimality condition for problem (1) with Banach space valued constraints, namely the generalized FJ optimality condition, which is customized for the application to the C-(E)MFC problem in Section 3.
We first state the generalized FJ optimality condition for problem (1).
Theorem 2.7.
Let $\widehat{x}\in C$ solve problem (1) locally. Then, there exist some $r_{0}\in\mathbb{R}$, $\lambda_{i}\in\mathtt{Y}_{i}^{*}$ for $i\in I$, and $\mu_{j}\in\mathtt{H}_{j}$ for $j\in J$, not all zero, such that
  • (i)
    (dual feasible) $r_{0}\geq 0$ and $\lambda_{i}\in K_{i}^{+}$ for all $i\in I$,
  • (ii)
    (complementary slackness) $\langle g_{i}(\widehat{x}),\lambda_{i}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}=0$, for all $i\in I$,
  • (iii)
    (non-trivial multipliers) $|r_{0}|+\sum_{i\in I}\lVert\lambda_{i}\rVert_{\mathtt{Y}_{i}^{*}}+\sum_{j\in J}\lVert\mu_{j}\rVert_{\mathtt{H}_{j}}=1$, where $\lVert\cdot\rVert_{\mathtt{H}_{j}}$ is the norm induced by the inner product $\langle\cdot,\cdot\rangle_{\mathtt{H}_{j}}$,
  • (iv)
    (Fritz-John condition) there exists $\xi\in N_{C}(\widehat{x})$ such that
    $-\xi=r_{0}Df(\widehat{x})+\sum_{i\in I}\lambda_{i}\circ Dg_{i}(\widehat{x})-\sum_{j\in J}\mu_{j}\circ Dh_{j}(\widehat{x}),$ (6)
    where $\lambda_{i}\circ Dg_{i}(\widehat{x})\in\mathtt{X}^{*}$ is the composition of $\lambda_{i}\in\mathtt{Y}_{i}^{*}$ and $Dg_{i}(\widehat{x})\in\mathcal{L}(\mathtt{X};\mathtt{Y}_{i})$. Similar notation holds for $\mu_{j}\circ Dh_{j}(\widehat{x})\in\mathtt{X}^{*}$.
Remark 2.8.
When $\widehat{x}\in C^{\rm o}$ in Theorem 2.7, we have $N_{C}(\widehat{x})=\{0\}$, which yields that the element $\xi$ stated in Theorem 2.7-(iv) can only be zero.
Before proving Theorem 2.7, let us first recall the Ekeland’s variation principle as below.
Lemma 2.9 (Ekeland’s variation principle [23]).
Let $(V,d)$ be a complete metric space and $F:C\subset V\mapsto\mathbb{R}$ be a lower semi-continuous function that is bounded from below. Assume that there exist some $\widehat{x}\in C$ and $\epsilon>0$ such that $F(\widehat{x})\leq\inf_{v\in C}F(v)+\epsilon$. Then, there exists a $z_{\epsilon}\in C$ such that (i) $F(z_{\epsilon})\leq F(\widehat{x})$, (ii) $d(\widehat{x},z_{\epsilon})\leq\sqrt{\epsilon}$, (iii) $F(v)+\sqrt{\epsilon}d(v,z_{\epsilon})\geq F(z_{\epsilon})$ for all $v\in C$.
Proof of Theorem 2.7.
Let $\epsilon\in(0,1)$ and recall $F:\mathtt{X}\to\mathbb{R}$ given in Lemma 2.9 that, for $x\in C$,
$F(x):=\max\left\{f(x)-f(\widehat{x})+\epsilon,~{}\!\!\!\!\sup_{\begin{subarray}{c}y_{i}^{*}\in K_{i}^{+}\\ \lVert y_{i}^{*}\rVert_{\mathtt{Y}_{i}^{*}}=1\end{subarray}}\langle g_{i}(x),y_{i}^{*}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}},~{}\lVert h_{j}(x)\rVert_{\mathtt{H}_{j}};~{}i\in I,~{}j\in J\right\}.$
Then, $x\mapsto F(x)$ is locally Lipschitz and bounded below by 00. As $\widehat{x}$ solves the problem (1) locally, we have
$F(\widehat{x})$ $=\max\left\{\epsilon,~{}\!\!\!\!\sup_{\begin{subarray}{c}y_{i}^{*}\in K_{i}^{+}\\ \lVert y_{i}^{*}\rVert_{\mathtt{Y}_{i}^{*}}=1\end{subarray}}\langle g_{i}(\widehat{x}),y_{i}^{*}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}},~{}0;~{}i\in I,~{}j\in J\right\}=\epsilon,$
and $\widehat{x}$ is an $\epsilon$-minimum of $x\mapsto F(x)$ over $C$. Thus, we can apply Lemma 2.9 to conclude the existence of $z_{\epsilon}\in C$ such that the assertions in Lemma 2.9 hold.
Note that $F(z_{\epsilon})>0$, otherwise $z_{\epsilon}\in C$ would be a feasible point due to Lemma 2.4, which contradicts the fact that $\widehat{x}$ solves problem (1) locally. By the statement under Remark 2.3, one can deduce that $z_{\epsilon}$ is a local minimum for the mapping $x\mapsto F(x)+\sqrt{\epsilon}\lVert x-z_{\epsilon}\rVert_{\mathtt{X}}+c_{F}d_{C}(x)$ with $c_{F}$ being the local Lipschitz constant of the mapping $F$ near $\widehat{x}$.
In lieu of Lemma 2.1 and Lemma 2.2, there exists a convex combination $(r_{0}^{\epsilon},\phi_{i}^{\epsilon},\psi_{j}^{\epsilon})$ for $i\in I$ and $j\in J$ such that the set
$r_{0}^{\epsilon}Df(z_{\epsilon})+\sum_{i\in I}\phi_{i}^{\epsilon}\partial\sup_{\begin{subarray}{c}y_{i}^{*}\in K_{i}^{+}\\ \lVert y_{i}^{*}\rVert_{\mathtt{Y}_{i}^{*}}=1\end{subarray}}\langle g_{i}(z_{\epsilon}),y_{i}^{*}\rangle_{{\mathtt{Y}_{i}},{\mathtt{Y}_{i}}^{*}}-\sum_{j\in J}\psi_{j}^{\epsilon}\partial\lVert h_{j}(z_{\epsilon})\rVert_{\mathtt{H}_{j}}+c_{F}\partial d_{C}(z_{\epsilon})$
intersects $\sqrt{\epsilon}B_{1}^{*}$ where $B_{1}^{*}$ is the unit ball in $\mathtt{X}^{*}$, and $\phi_{i}^{\epsilon}=0$ if $g_{i}(z_{\epsilon})\leq_{K_{i}}0$ for $i\in I$ and $\psi_{j}^{\epsilon}=0$ if $h_{j}(z_{\epsilon})=0$ for $j\in J$. It is easy to verify that $x\mapsto\lVert h_{j}(x)\rVert_{\mathtt{H}_{j}}$ is continuously Fréchet differentiable and $D\lVert h_{j}(x)\rVert_{\mathtt{H}_{j}}=(h_{j}(x)/\lVert h_{j}(x)\rVert_{\mathtt{H}_{j}})\circ Dh_{j}(x)$ when $h_{j}(x)\neq 0$ for $i\in I$. Thus, let us set
$\mu_{j}^{\epsilon}:=-(h_{j}(z_{\epsilon})/\lVert h_{j}(z_{\epsilon})\rVert_{\mathtt{H}_{j}})\psi_{j}^{\epsilon}\in\mathtt{H}_{j}$
when $h_{j}(z_{\epsilon})=0$, and set $\mu_{j}^{\epsilon}:=0$ otherwise, for $j\in J$.
On the other hand, we have
$D\langle g_{i}(x),y_{i}^{*}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}=y_{i}^{*}\circ Dg_{i}(x)\in\mathtt{X}^{*}$
for $x\in\mathtt{X}$. We claim that the set $\{y_{i}^{*}\circ Dg_{i}(x);~{}y_{i}^{*}\in M(x)\}$ is closed and convex, where $M_{i}(x)$ is the maximizer set given by
$M_{i}(x)=\left\{y_{i}^{*}\in\{y_{i}^{*}\in K_{i}^{+};~{}\lVert y_{i}^{*}\rVert_{\mathtt{Y}_{i}^{*}}=1\};~{}\langle g_{i}(x),y_{i}^{*}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}=\sup_{\begin{subarray}{c}y_{i}^{*}\in K_{i}^{+}\\ \lVert y_{i}\rVert_{\mathtt{Y}_{i}^{*}}=1\end{subarray}}\langle g_{i}(x),y_{i}^{*}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\right\}.$
As $\langle g_{i}(x),y_{i}^{*}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}$ is weak-* continuous in $y_{i}^{*}$ for any $x\in\mathtt{X}$ and $\{y_{i}^{*}\in K_{i}^{+};~{}\lVert y_{i}^{*}\rVert_{\mathtt{Y}_{i}^{*}}=1\}$ is weak-* compact due to Alaogu’s theorem, $M_{i}(x)$ is nonempty, and moreover it can be easily verified to be convex and weak-* closed (and hence closed). The convexity of $\{y_{i}^{*}\circ Dg_{i}(x);~{}y_{i}^{*}\in M_{i}(x)\}$ follows from $Dg_{i}(x)\in\mathcal{L}(\mathtt{X},\mathtt{Y}_{i})$ and the convexity of $M_{i}(x)$. To show the closedness, let us assume that $y_{i}^{n^{*}}\circ Dg_{i}(x)\to x^{*}$ in $\mathtt{X}^{*}$ with $y_{i}^{n^{*}}\in M_{i}(x)$. As the latter set is bounded, according to Lemma 2.5-(i), there exists a subsequence $(y_{i}^{n_{k}^{*}})$ of $(y_{i}^{n^{*}})$ , weakly- converging to some $y^{*}\in\mathtt{Y}_{i}^{*}$ (in fact $y^{*}\in M_{i}(x)$ in view that $M_{i}(x)$ is weak- closed). Thus, we can conclude that
$y^{*}\circ Dg_{i}(x)(x^{\prime})=\lim_{n_{k}\to\infty}y_{i}^{n_{k}^{*}}\circ Dg_{i}(x)(x^{\prime})=\langle x^{\prime},x^{*}\rangle_{\mathtt{X},\mathtt{X}^{*}},\quad\forall x^{\prime}\in\mathtt{X}.$
Hence, we have $x^{*}=y^{*}\circ Dg_{i}(x)$. It then follows from Lemma 2.2 again that, for $i\in I$,
$\partial\sup_{\begin{subarray}{c}y_{i}^{*}\in K_{i}^{+}\\ \lVert y_{i}\rVert_{\mathtt{Y}_{i}^{*}}=1\end{subarray}}\langle g_{i}(x),y_{i}^{*}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}$ $\subset\overline{\mathrm{co}\left\{y_{i}^{*}\circ Dg_{i}(x);~{}y_{i}\in M_{i}(x)\right\}}=\left\{y_{i}^{*}\circ Dg_{i}(x);~{}y_{i}\in M_{i}(x)\right\}.$
We can then deduce that the element in $\partial\sup_{{y_{i}^{*}\in K_{i}^{+},~{}\lVert y_{i}\rVert_{\mathtt{Y}_{i}^{*}}=1}}\langle g_{i}(x),y_{i}^{*}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}$ can be expressed in the form of $y_{i}^{*}\circ Dg_{i}(x)$ with $y_{i}^{*}\in M_{i}(x)$, and thus $\|y_{i}^{*}\|_{\mathtt{Y}_{i}^{*}}=1$. We then set $\lambda_{i}^{\epsilon}=\phi_{i}^{\epsilon}y_{i}^{*}\in K_{i}^{+}$ with $y_{i}^{*}$ given above for $i\in I$. Recall $\phi_{i}^{\epsilon}=0$ when $g_{i}(z_{\epsilon})\leq_{K_{i}}0$, and it thus holds that $\exists~{}\xi_{\epsilon}\in\mathtt{X}^{*}$ such that
$\begin{cases}\displaystyle-\xi_{\epsilon}=r_{0}^{\epsilon}Df(z_{\epsilon})+\sum_{i\in I}\lambda_{i}^{\epsilon}\circ Dg_{i}(z_{\epsilon})-\sum_{j\in J}\mu_{j}^{\epsilon}\circ Dh_{j}(z_{\epsilon}),\\[8.00003pt] \displaystyle~{}~{}\xi_{\epsilon}\in c_{F}\partial d_{C}(z_{\epsilon})-\sqrt{\epsilon}B_{1}^{*},\\[6.00006pt] \displaystyle~{}~{}r_{0}^{\epsilon}\geq 0,~{}\lambda_{i}^{\epsilon}\in K_{i}^{+},~{}\lambda_{i}^{\epsilon}=0\text{ if }g_{i}(z_{\epsilon})\leq_{K_{i}}0,~{}\forall i\in I,\\[8.00003pt] \displaystyle\langle g_{i}(z_{\epsilon}),\lambda_{i}^{\epsilon}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}=\phi_{i}^{\epsilon}\sup_{\begin{subarray}{c}y_{i}^{*}\in K_{i}^{+}\\ \lVert y_{i}^{*}\rVert_{\mathtt{Y}_{i}^{*}}=1\end{subarray}}\langle g_{i}(z_{\epsilon}),y_{i}^{*}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\leq F(z_{\epsilon})\leq\epsilon,~{}\forall i\in I,\\[18.00005pt] \displaystyle~{}~{}r_{0}^{\epsilon}+\sum_{i\in I}\|\lambda_{i}^{\epsilon}\|_{\mathtt{Y}_{i}^{*}}+\sum_{j\in J}\lVert\mu_{j}^{\epsilon}\rVert_{\mathtt{H}_{j}}=1.\end{cases}$
The last equality results from the fact that $(r_{0}^{\epsilon},\phi_{i}^{\epsilon},\psi_{j}^{\epsilon})$ for $i\in I$ and $j\in J$ is a convex combination and they multiply functionals with unit norms. Now, let $\epsilon\to 0$, by the definition of $z_{\epsilon}\in C$, we have $z_{\epsilon}\to\widehat{x}$. In view of the boundedness, we may assume that $(\xi_{\epsilon},\lambda_{i}^{\epsilon},\mu_{j}^{\epsilon})$ converges to $(\xi,\lambda_{i},\mu_{j})$ in the weak-* topology and $r_{0}^{\epsilon}$ converges to $r_{0}\geq 0$ (by a subsequence if necessary) as $\epsilon\to 0$. On the other hand, Lemma 2.1-(v) implies both (6) and the relation $\xi\in c_{F}\partial d_{C}(\widehat{x})\subset N_{C}(\widehat{x})$. In addition, $\lambda_{i}\in K_{i}^{+}$ as $K_{i}^{+}$ is closed and convex for $i\in I$. Obviously, Theorem 2.7-(iii) holds and it suffices to show (ii). For any $i\in I$, if $g_{i}(z_{\epsilon})\leq_{K_{i}}0$, then $\langle g_{i}(z_{\epsilon}),\lambda_{i}^{\epsilon}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}=0$, and if $-g_{i}(z_{\epsilon})\notin K_{i}$, then Lemma 2.4 results in $\langle g_{i}(z_{\epsilon}),\lambda_{i}^{\epsilon}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\geq 0$. All in all, we have $|\langle g_{i}(z_{\epsilon}),\lambda_{i}^{\epsilon}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}|\leq\epsilon$. Moreover, we can prove $\langle g_{i}(\widehat{x}),\lambda_{i}\rangle_{Y_{i},Y_{i}^{*}}=\lim_{\epsilon\to 0}|\langle g_{i}(z_{\epsilon}),\lambda_{i}^{\epsilon}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}|=0$ as in Lemma 2.6, which yields Theorem 2.7-(ii). In particular, Theorem 2.7-(iii) also implies that the multipliers are not all zeros, which completes the proof. ∎
Remark 2.10.
Thanks to Lemma 2.6, it follows from Theorem 2.7-(ii) that $\lambda_{i}=0$ when $g_{i}(\widehat{x})<_{K_{i}}0$.
2.3 Constraint Qualifications
The goal of this subsection is to discuss the constraint qualification (CQ) conditions under which our generalized Fritz-John optimality condition can be enhanced to be the Karush-Kuhn-Tucker (KKT) type condition, which plays an important role in studying the classical constrained stochastic control problems in Section 4.
Let us assume that the space $\mathtt{X}$ admits the decomposition $\mathtt{X}=\mathtt{X}_{1}\times\mathtt{X}_{2}$, where both $(\mathtt{X}_{1},\|\cdot\|_{\mathtt{X}_{1}})$ and $(\mathtt{X}_{2},\|\cdot\|_{\mathtt{X}_{2}})$ are Banach spaces. The norm on $\mathtt{X}$ is defined to be the product norm $\|(x_{1},x_{2})\|_{\mathtt{X}}=\|x_{1}\|_{\mathtt{X}_{1}}+\|x_{2}\|_{\mathtt{X}_{2}}$ for $(x_{1},x_{2})\in\mathtt{X}_{1}\times\mathtt{X}_{2}$. We also define the natural projection mapping $\Pi_{1}:\mathtt{X}\to\mathtt{X}_{1}$ which maps $x=(x_{1},x_{2})\in\mathtt{X}=\mathtt{X}_{1}\times\mathtt{X}_{2}$ to $x_{1}\in\mathtt{X}_{1}$. Then, we introduce the following definitions:
Definition 2.11.
Let $\widehat{x}=(\widehat{x}_{1},\widehat{x}_{2})\in C$ be feasible for problem (1), and set $I(\widehat{x}):=\{i\in I;~{}g_{i}(\widehat{x})\in\partial K_{i}\}$. We say that
  • (i)
    the linear independence constraint qualification (LICQ) holds at $\widehat{x}$ (write ${\rm LICQ}(\widehat{x})$) if the following mapping is surjective:
    $(D_{x_{1}}g_{i}(\widehat{x}_{1}),D_{x_{1}}h_{j}(\widehat{x}_{1}))_{i\in I(\widehat{x}),j\in J}:\mathtt{X}_{1}\mapsto\prod_{i\in I(\widehat{x})}\mathtt{Y}_{i}\times\prod_{j\in J}\mathtt{H}_{j};$
  • (ii)
    the Mangasarian-Fromovitz constraint qualification (MFCQ) holds at $\widehat{x}$ (write ${\rm MFCQ}(\widehat{x})$) if the mapping $(D_{x_{1}}h_{j}(\widehat{x}_{1}))_{j\in J}:\mathtt{X}_{1}\mapsto\prod_{j\in J}\mathtt{H}_{j}$ is surjective, and there exists an element $c\in C$ such that
    $Dg_{i}(\widehat{x})(c-\widehat{x})<_{K_{i}}0,~{}\forall i\in I(\widehat{x}),\quad Dh_{j}(\widehat{x})(c-\widehat{x})=0,~{}\forall j\in J.$
Remark 2.12.
If $\mathtt{X}$ is a Euclidean space, $\mathtt{X}_{1}=\mathtt{X}$ and $\mathtt{X}_{2}=\{0\}$, the CQ conditions defined in Definition 2.11 reduce to LICQ and MFCQ in the Euclidean space, respectively.
Next, we also provide a sufficient condition for the KKT optimality condition to hold. Let $\Pi_{1}(C):=\{\Pi_{1}(x);~{}x\in C\}$. Then
Corollary 2.13.
Let $\widehat{x}\in C$ solve problem (1) locally and $\Pi_{1}(\widehat{x})\in(\Pi_{1}(C))^{\rm o}$ in the relative topology of $\mathtt{X}_{1}$. If either ${\rm LICQ}(\widehat{x})$ or ${\rm MFCQ}(\widehat{x})$ is fulfilled, the KKT optimality condition holds true in the sense that there exist $\lambda_{i}\in\mathtt{Y}_{i}^{*}$ and $\mu_{j}\in\mathtt{H}_{j}$ for $i\in I$ and $j\in J$ such that
  • (i)
    (dual feasible) $\lambda_{i}\in K_{i}^{+}$ for all $i\in I$,
  • (ii)
    (complementary slackness) $\langle g_{i}(\hat{x}),\lambda_{i}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}=0$ for all $i\in I$,
  • (iii)
    (KKT condition) there exists $\xi\in N_{C}(\widehat{x})$ such that
    $-\xi=Df(\widehat{x})+\sum_{i\in I}\lambda_{i}\circ Dg_{i}(\widehat{x})-\sum_{j\in J}\mu_{j}\circ Dh_{j}(\widehat{x}).$
Proof.
We first assume that MFCQ($\widehat{x}$) holds. If Corollary 2.13-(iii) does not hold, then $r_{0}=0$, we can deduce that there exists a $\xi\in N_{C}(\widehat{x})$ such that
$-\xi=\sum_{i\in I}\lambda_{i}\circ Dg_{i}(\widehat{x})-\sum_{j\in J}\mu_{j}\circ Dh_{j}(\widehat{x}),$ (7)
where $(\lambda_{i})_{i\in I}$ and $(\mu_{j})_{j\in J}$ can not be zero simultaneously thanks to Theorem 2.7-(iv). By the definition of normal cone $N_{C}(\widehat{x})$, MFCQ($\widehat{x}$) and Lemma 2.6, the equality (7) can be written as
$0\leq\langle c-\widehat{x},\xi\rangle=\sum_{i\in I}\langle Dg_{i}(\widehat{x})(c-\widehat{x}),\lambda_{i}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}-\sum_{j\in J}\langle Dh_{j}(\widehat{x})(c-\widehat{x}),\mu_{j}\rangle_{\mathtt{H}_{j}}\leq 0$
In above, the equality can only be attained when $\lambda_{i}=0$ for $i\in I$. Thus, the equality (7) can be translated into
$\sum_{j\in J}\mu_{j}\circ Dh_{j}(\widehat{x})=\xi\in N_{C}(\widehat{x}).$
Here, $\xi=(\xi_{1},\xi_{2})\in\mathtt{X}_{1}^{*}\times\mathtt{X}_{2}^{*}$ as the dual space of a product space is the product of their dual spaces. Hence, we have $\sum_{j\in J}\mu_{j}\circ D_{x_{1}}h_{j}(\widehat{x})=\xi_{1}$. Moreover, by the definition of normal cone, we can conclude that by choosing $c=(c_{1},\widehat{x}_{2})\in C\subset\mathtt{X}_{1}\times\mathtt{X}_{2}$, where we recall $\widehat{x}=(\widehat{x}_{1},\widehat{x}_{2})\in C\subset\mathtt{X}$, we have
$0\leq\langle\widehat{x}-c,\xi\rangle_{\mathtt{X},\mathtt{X}^{*}}=\langle\widehat{x}_{1}-c_{1},\xi_{1}\rangle_{\mathtt{X}_{1},\mathtt{X}_{1}^{*}}.$ (8)
As $\widehat{x}_{1}=\Pi_{1}(\widehat{x})\in(\Pi_{1}(C))^{\rm o}$, we deduce that
$\sum_{j\in J}\mu_{j}\circ D_{x_{1}}h_{j}(\widehat{x}_{1})=\xi_{1}=0.$ (9)
Due to the fact that $(D_{x_{1}}h_{j}(\widehat{x}_{1}))_{j\in J}:\mathtt{X}_{1}\to\prod_{j\in J}\mathtt{H}_{j}$ is surjective, we get that $\mu_{j}=0$ $(j\in J)$, which contradicts the assertion in Lemma 2.7-(iii). Thus, Corollary 2.13-(iii) holds.
Now, let us assume that LICQ($\widehat{x}$) holds. From (7) and (8) again, it follows that
$\sum_{i\in I}\lambda_{i}\circ D_{x_{1}}g_{i}(\widehat{x}_{1})-\sum_{j\in J}\mu_{j}\circ D_{x_{1}}h_{j}(\widehat{x}_{1})=-\xi_{1}=0.$
We argue by contradiction to show that $\lambda_{i}=0$ for $i\in I$. To this end, we first suppose that $\ker(\lambda_{1})^{c}\neq\varnothing$. Then, by the surjection introduced in LICQ($\widehat{x}$), there exists an $x_{0}\in\mathtt{X}_{1}$ such that $(D_{x_{1}}g_{i}(\widehat{x}_{1})(x_{0}),D_{x_{1}}h_{j}(\widehat{x}_{1})(x_{1}))_{i\in I,j\in J}$ belongs to the nonempty set $\ker(\lambda_{1})^{c}\times\prod_{i\in I\backslash\{1\}}\ker(\lambda_{i})\times\prod_{j\in J}\ker(\mu_{j})$, which yields that
$0=\sum_{i\in I}\lambda_{i}\circ D_{x_{1}}g_{i}(\widehat{x}_{1})(x_{0})-\sum_{j\in J}\mu_{j}\circ D_{x_{1}}h_{j}(\widehat{x}_{1})(x_{0})=\left\langle D_{x_{1}}g_{1}(\widehat{x}_{1})(x_{0}),\lambda_{1}\right\rangle_{\mathtt{Y}_{1},\mathtt{Y}_{1}^{*}}\neq 0$
This contradiction leads to $\ker(\lambda_{1})=\mathtt{Y}_{1}$, and hence $\lambda_{1}=0$. We then arrive at
$\sum_{i\in I\backslash\{1\}}\lambda_{i}\circ D_{x_{1}}g_{i}(\widehat{x}_{1})-\sum_{j\in J}\mu_{j}\circ D_{x_{1}}h_{j}(\widehat{x}_{1})=0.$
Repeating the procedures above, we conclude that $\lambda_{i}=0$ for all $i\in I$. Hence, the above equality reduces to (9). Following the proof in the MFCQ case, we can deduce that $\mu_{j}=0$ for $j\in J$ as desired. ∎
3 Extended MFC Problems with Constraints
In this section, we apply the newly obtained FJ optimality condition in Theorem 2.7 to establish the SMP for some C-(E)MFC problems in which we can encompass both dynamic expectations constraints and mixed state-control constraints. To this end, we first formulate our C-(E)MFC problem and introduce some basic spaces and notations as preparations.
Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered probability space with the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ satisfying the usual conditions. For $r,l\in\mathbb{N}$, let $(W_{t})_{t\in[0,T]}$ be a standard $r$-dimensional $(\mathbb{P},\mathbb{F})$-Brownian motion and denote $U\subset\mathbb{R}^{l}$ as the closed and convex control space.
We define $\mathcal{X}$ as the space $L^{2}([0,T];\mathbb{R}^{l})/\mathord{\sim}$ equipped with the inner product $\langle\gamma^{1},\gamma^{2}\rangle_{\mathcal{X}}=\int_{0}^{T}\langle\gamma_{t}^{1},\gamma_{t}^{2}\rangle\mathrm{d}t$. Then, $(\mathcal{X},\langle\cdot,\cdot\rangle_{\mathcal{X}})$ is a Hilbert space. For any $\gamma=(\gamma_{t})_{t\in[0,T]}\in\mathcal{X}$, we introduce, for fixed $t\in[0,T]$,
$\widetilde{\gamma}^{t}_{s}:=\gamma_{s}\mathbf{1}_{s\in[0,t]},\quad\forall s\in[0,T].$ (10)
Then $\|\widetilde{\gamma}^{t}\|_{\mathcal{X}}=\sqrt{\langle\widetilde{\gamma}^{t},\widetilde{\gamma}^{t}\rangle_{\mathcal{X}}}\leq\|\gamma\|_{\mathcal{X}}$ for all $t\in[0,T]$. Denote by $\mathscr{U}[0,T]$ the set of admissible controls $\alpha=(\alpha_{t})_{t\in[0,T]}$ valued in $U$ such that the process $\alpha$ is $\mathbb{F}$-adapted and satisfies $\mathbb{E}[\int_{0}^{T}|\alpha_{t}|^{2}\mathrm{d}t]<\infty$. Let $\mathbb{L}_{\mathbb{F}}^{2}$ be the set of $\mathbb{F}$-adapted processes $\beta=(\beta_{t})_{t\in[0,T]}$ taking values in $\mathbb{R}^{l}$ such that $\mathbb{E}[\int_{0}^{T}|\beta_{t}|^{2}\mathrm{d}t]<\infty$. Then, $\mathbb{L}_{\mathbb{F}}^{2}$ is a Hilbert space equipped with the inner product that
$\langle\beta^{1},\beta^{2}\rangle_{\mathbb{L}_{\mathbb{F}}^{2}}=\mathbb{E}\left[\int_{0}^{T}\langle\beta^{1}_{t},\beta^{2}_{t}\rangle\mathrm{d}t\right],\quad\forall\beta^{1},\beta^{2}\in\mathbb{L}_{\mathbb{F}}^{2},$ (11)
where $\beta^{1}=\beta^{2}$ means that $\beta^{1}(\omega)=\beta^{2}(\omega)$ holds in $\mathcal{X}$ except for a $\mathbb{P}$-null set. It can be easily verified that $\mathscr{U}[0,T]$ is closed and convex in $(\mathbb{L}_{\mathbb{F}}^{2},\langle\cdot,\cdot\rangle_{\mathbb{L}_{\mathbb{F}}^{2}})$. We can observe that $\beta\in L^{2}(\Omega;\mathcal{X})$ for all $\beta\in\mathbb{L}_{\mathbb{F}}^{2}$, and moreover, we have
$\|\beta\|_{\mathbb{L}_{\mathbb{F}}^{2}}^{2}:=\mathbb{E}\left[\|\beta\|_{\mathcal{X}}^{2}\right]=\mathbb{E}\left[\langle\beta,\beta\rangle_{\mathcal{X}}\right].$ (12)
Let us define $\mathbb{M}^{2}$ ($\subset\mathbb{L}_{\mathbb{F}}^{2}$) as the set of $\mathbb{F}$-adapted (continuous) $\mathbb{R}^{n}$-valued process $\eta=(\eta_{t})_{t\in[0,T]}$ satisfying the Itô’s representation that
$\eta_{t}=\eta_{0}+\int_{0}^{t}\eta^{1}_{s}\mathrm{d}s+\int_{0}^{t}\eta_{s}^{2}\mathrm{d}W_{s},$ (13)
where $\eta_{0}\in L^{2}(\Omega;\mathbb{R}^{n})$ is $\mathcal{F}_{0}$-measurable, $\eta^{1}=(\eta_{t}^{1})_{t\in[0,T]}$ and $\eta^{2}=(\eta_{t}^{2})_{t\in[0,T]}$ take values respectively in $\mathbb{R}^{n}$ and $\mathbb{R}^{n\times r}$, which are $\mathbb{F}$-adapted processes such that $\mathbb{E}[\int_{0}^{T}(|\eta_{t}^{1}|^{2}+|\eta_{t}^{2}|^{2})\mathrm{d}t]<\infty$. Then, $\mathbb{M}^{2}$ becomes a Hilbert space under the inner product $\langle\cdot,\cdot\rangle_{\mathbb{M}^{2}}$ defined by
$\langle\eta,\zeta\rangle_{\mathbb{M}^{2}}:=\mathbb{E}\left[\langle\eta_{0},\zeta_{0}\rangle+\int_{0}^{T}\left(\langle\eta_{t}^{1},\zeta_{t}^{1}\rangle+\mathrm{tr}\left(\eta_{t}^{2}(\zeta_{t}^{2})^{\top}\right)\right)\mathrm{d}t\right],\quad\forall\eta,\zeta\in\mathbb{M}^{2}.$ (14)
We use $\mathbb{L}_{\mathbb{F},\mathbb{R}}^{2}$ to represent the set of adapted processes in $L^{2}([0,T]\times\Omega;\mathbb{R})/\mathord{\sim}$. Then, $\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}$ becomes a Hilbert space when it is equipped with the inner product $\langle\cdot,\cdot\rangle_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}$ defined by
$\langle\beta^{1},\beta^{2}\rangle_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}:=\mathbb{E}\left[\int_{0}^{T}\beta_{t}^{1}\beta_{t}^{2}\mathrm{d}t\right],\quad\forall\beta^{1},\beta^{2}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}.$ (15)
We also denote $\|\cdot\|_{\mathbb{L}_{\mathbb{F},\mathbb{R}}^{2}}$ as the induced norm by above inner product.
3.1 Partial L-derivative
As a technical preparation, we follow Acciaio et al. [1] to introduce the partial L-derivative of joint probability laws (i.e. probability measures on product spaces).
Let $w:\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\mapsto\mathbb{R}$. We use $\rho$ to represent a generic element in $\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})$, while $\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})$ and $\nu\in\mathcal{P}_{2}(\mathbb{R}^{l})$ denote its respective marginals. For a probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}})$, define the lift $\tilde{w}:L^{2}((\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}});\mathbb{R}^{n}\times\mathbb{R}^{l})$ of the element $w$ by
$\tilde{w}(\tilde{X},\tilde{\alpha})=w\left(\mathcal{L}(\tilde{X},\tilde{\alpha})\right),$ (16)
where $\mathcal{L}(\tilde{X},\tilde{\alpha})$ denotes the joint distribution of $(\tilde{X},\tilde{\alpha})$ under $\tilde{P}$. Then, the L-derivative is defined by
Definition 3.1.
We call $w$ the L-differentiable if there is a pair $(\tilde{X},\tilde{\alpha})\in L^{2}((\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}});\mathbb{R}^{n}\times\mathbb{R}^{l})$ with the joint distribution $\mathcal{L}^{\tilde{P}}(\tilde{X},\tilde{\alpha})$ such that the lifted function $\tilde{w}$ is Fréchet differentiable at $(\tilde{X},\tilde{\alpha})$.
By Section 5.2.1 in Carmona and Delarue [15], there exists a measurable mapping $\partial_{\rho}w(\rho)(\cdot):\mathbb{R}^{n}\times\mathbb{R}^{l}\mapsto\mathbb{R}^{n}\times\mathbb{R}^{l}$ such that $\partial_{\rho}w(\rho)(\tilde{X},\tilde{\alpha})=D\tilde{w}(\tilde{X},\tilde{\alpha})$. Since $L^{2}((\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}});\mathbb{R}^{n}\times\mathbb{R}^{l})\simeq L^{2}((\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}});\mathbb{R}^{n})\times L^{2}((\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}});\mathbb{R}^{l})$, we can define the first component and second component of $\partial_{\rho}w(\rho)(\tilde{X},\tilde{\alpha})$ as $\partial_{\mu}w(\rho)(\tilde{X},\tilde{\alpha})$ and $\partial_{\nu}w(\rho)(\tilde{X},\tilde{\alpha})$, respectively. By the measurability of $\partial_{\rho}w(\rho)(\tilde{X},\tilde{\alpha})$, we can also conclude that $\partial_{\mu}w(\rho)(\cdot):\mathbb{R}^{n}\times\mathbb{R}^{l}\mapsto\mathbb{R}^{n}$ and $\partial_{\nu}w(\rho)(\cdot):\mathbb{R}^{n}\times\mathbb{R}^{l}\mapsto\mathbb{R}^{l}$ are both measurable. We call $\partial_{\mu}w(\rho)$ and $\partial_{\nu}w(\rho)$ the partial derivatives of $w$. By Definition 3.1 on the L-derivative, we have the following lemma.
Lemma 3.2.
Let $\rho=\mathcal{L}(\tilde{X},\tilde{\alpha})$ for a $\mathbb{R}^{n}\times\mathbb{R}^{l}$-valued r.v. $(\tilde{X},\tilde{\alpha})$ on some probability space $(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{P})$. Then, for any $\mathbb{R}^{n}\times\mathbb{R}^{l}$-valued r.v. $(\tilde{X}^{\prime},\tilde{\alpha}^{\prime})$, it holds that
$w(\mathcal{L}(\tilde{X}^{\prime},\tilde{\alpha}^{\prime}))$ $=w(\rho)+\mathbb{E}^{\tilde{P}}\left[\langle\partial_{\mu}(\rho)(\tilde{X},\tilde{\alpha}),\tilde{X}^{\prime}-\tilde{X}\rangle\right]+\mathbb{E}^{\tilde{P}}\left[\langle\partial_{\nu}(\rho)(\tilde{X},\tilde{\alpha}),\tilde{\alpha}^{\prime}-\tilde{\alpha}\rangle\right]$
$\quad+o\left(\sqrt{\mathbb{E}^{\tilde{P}}[|(\tilde{X}^{\prime},\tilde{\alpha}^{\prime})-(\tilde{X},\tilde{\alpha})|^{2}]}\right),$
[…]
In the sequel, we will construct a copy space $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{P}^{\prime})$ of the probability space $(\Omega,\mathcal{F},\mathbb{P})$ in the sense that $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{P}^{\prime})=(\Omega,\mathcal{F},\mathbb{P})$. For a random variable $X$ constructed on $(\Omega,\mathcal{F},\mathbb{P})$, we denote by $X^{\prime}$ its copy random variable on $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{P}^{\prime})$ such that $X^{\prime}(\omega^{\prime})=X(\omega)$ for $\omega^{\prime}=\omega\in\Omega$.
3.2 Formulation of C-(E)MFC problems
In this section, we formulate our C-(E)MFC problems. Let $(b,\sigma):[0,T]\times\mathbb{R}^{n}\times{\cal P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathbb{R}^{l}\mapsto\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ be measurable mappings. For any $\alpha=(\alpha_{t})_{t\in[0,T]}\in\mathscr{U}[0,T]$, consider the following controlled Mckean-Vlasov SDE that
$\mathrm{d}X_{t}^{\alpha}$ $=b(t,X_{t}^{\alpha},\mathcal{L}(X_{t}^{\alpha},\alpha_{t}),\alpha_{t})\mathrm{d}t+\sigma(t,X_{t}^{\alpha},\mathcal{L}(X_{t}^{\alpha},\alpha_{t}),\alpha_{t})\mathrm{d}W_{t},$
$X_{0}^{\alpha}$ $=\kappa\in L^{2}((\Omega,\mathcal{F}_{0},\mathbb{P});\mathbb{R}^{n}),$ (17)
where $\mathcal{L}(X_{t}^{\alpha},\alpha_{t})$ denotes the joint distribution of $(X_{t}^{\alpha},\alpha_{t})$ for $t\in[0,T]$.
Let integer $1\leq m_{1}\leq m$. For $\alpha\in\mathscr{U}[0,T]$ and $\phi^{i}:[0,T]\times\mathbb{R}^{n}\times\mathcal{X}\mapsto\mathbb{R}$ with $i\in I_{1}:=\{1,2,\ldots,m_{1}\}$, we introduce the dynamic expectation constraints in the sense that, for $i\in I_{1}$,
$\mathbb{E}\left[\phi^{i}\left(t,X_{t}^{\alpha},\widetilde{\alpha}^{t}\right)\right]\geq 0,\quad\forall t\in[0,T].$ (18)
For $\psi^{j}:[0,T]\times C([0,T];\mathbb{R}^{n})\times\mathbb{R}^{l}\mapsto\mathbb{R}$ with $j\in I_{2}:=I\setminus I_{1}$, the dynamic state-control constraint is formulated by, for $j\in I_{2}$,
$\psi^{j}\left(t,X_{t}^{\alpha},\alpha_{t}\right)\geq 0,\quad\mathrm{d}t\times\mathrm{d}\mathbb{P}\text{-}\mathrm{a.s.}.$ (19)
Here, $X^{\alpha}=(X_{t}^{\alpha})_{t\in[0,T]}$ obeys the controlled Mckean-Vlasov SDE (3.2) and the process $\widetilde{\alpha}^{t}$ is defined by (10). Then, we introduce the admissible control set $\mathscr{U}_{ad}[0,T]$ as the set of processes $\alpha\in\mathscr{U}[0,T]$ such that there exists $X^{\alpha}=(X_{t}^{\alpha})_{t\in[0,T]}\in\mathbb{M}^{2}$ satisfying (3.2), (18) and (19). Based on the preparations provided above, we introduce the cost functional as follows, for $\alpha=(\alpha_{t})_{t\in[0,T]}\in\mathscr{U}_{ad}[0,T]\neq\varnothing$,
$J(\alpha):=\mathbb{E}\left[\int_{0}^{T}f(t,X_{t}^{\alpha},\mathcal{L}(X_{t}^{\alpha},\alpha_{t}),\alpha_{t})\mathrm{d}t+g(X_{T}^{\alpha},\mathcal{L}(X_{T}^{\alpha}))\right].$ (20)
We make the following assumptions on the model coefficients throughout the section:
Assumption 1.
  • (A1)
    for $i\in I_{1}$ and $j\in I_{2}$, the mappings $b,\sigma,f,g,\phi^{i}$ and $\psi^{j}$ are Borel measurable and continuously differentiable w.r.t. $x\in\mathbb{R}^{n}$. Moreover, $b,\sigma,f,\psi^{j}$ are continuously differentiable w.r.t. $u\in\mathbb{R}^{l}$ and $\phi^{i}$ is continuously Frêchet differentiable w.r.t. $\gamma\in\mathcal{X}$ Moreover, $\phi^{i}$ is jointly continuous in $(t,x)\in[0,T]\times\mathbb{R}^{n}$.
  • (A2)
    there is a constant $M>0$ such that, for all $(t,x,u,\gamma)\in[0,T]\times\mathbb{R}^{n}\times\mathbb{R}^{l}\times\mathcal{X}$,
    $\sum_{i\in I_{1}}\left(|\nabla_{x}\phi^{i}(t,x,\gamma)|+\|D_{\gamma}\phi^{i}(t,x,\gamma)\|_{\mathcal{X}}\right)$ $\leq M\left(1+|x|+|u|+\|\gamma\|_{\mathcal{X}}\right),$
    $\sum_{j\in I_{2}}\left(|\nabla_{x}\psi^{j}(t,x,u)|+|\nabla_{u}\psi^{j}(t,x,u)|\right)$ $\leq M.$
  • (A3)
    the mappings $b(t,x,\rho,u)$ and $\sigma(t,x,\rho,u)$ are uniformly Lipschitz continuous in $(x,\rho)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})$ in the sense that, there is a constant $M>0$ independent of $u\in\mathbb{R}^{l}$ such that, for all $(x,\rho),(x^{\prime},\rho^{\prime})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})$ and $(t,u)\in[0,T]\times\mathbb{R}^{l}$,
    $\left|b(t,x^{\prime},\rho^{\prime},u)-b(t,x,\rho,u)\right|+\left|\sigma(t,x^{\prime},\rho^{\prime},u)-\sigma(t,x,\rho,u)\right|\leq M\left(|x-x^{\prime}|+\mathcal{W}_{2}(\rho,\rho^{\prime})\right),$
    where $\mathcal{W}_{2}(\cdot,\cdot)$ denotes the Wasserstein metric on ${\cal P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})$.
  • (A4)
    there is a constant $M>0$ such that, for all $(t,x,\rho,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathbb{R}^{l}$, it holds that
    $|\delta(t,x,\rho,u)|\leq M(1+|x|+|u|+M_{2}(\rho)),$
    where $\delta\in\{b,\sigma,\nabla_{u}b(\cdot),\nabla_{u}\sigma(\cdot),\nabla_{x}f(\cdot),\nabla_{u}f(\cdot)\}$ and $M_{2}(\rho):=\int_{\mathbb{R}^{n}\times\mathbb{R}^{l}}(|x|^{2}+|u|^{2})\rho(\mathrm{d}x,\mathrm{d}u)<\infty$. Moreover, for all $(x,\mu)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$,
    $|\nabla_{x}g(x,\mu)|\leq M(1+|x|+M_{2}(\mu)),$
    with $M_{2}(\mu):=\int_{\mathbb{R}^{n}}|x|^{2}\mu(\mathrm{d}x)<\infty$.
  • (A5)
    the mappings $b(t,x,\rho,u),\sigma(t,x,\rho,u),f(t,x,\rho,u)$ and $g(x,\mu)$ are continuously L-differentiable w.r.t. $\rho\in\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})$ or $\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})$. Moreover, there is a constant $M>0$ such that, for all $(t,x,\rho,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$,
    $\left(\int_{\mathbb{R}^{n}\times\mathbb{R}^{l}}|\partial_{\rho}f(t,x,\rho,u)(x^{\prime},u^{\prime})|^{2}\rho(\mathrm{d}x^{\prime},\mathrm{d}u^{\prime})\right)^{\frac{1}{2}}+\left(\int_{\mathbb{R}^{n}}|\partial_{\mu}g(x,\mu)(x^{\prime})|^{2}\mu(\mathrm{d}x^{\prime})\right)^{\frac{1}{2}}$
    $\qquad\qquad\leq M(1+|x|+|u|+M_{2}(\rho)+M_{2}(\mu)).$
We first have the following remark:
Remark 3.3.
[…]
$\left(\int_{\mathbb{R}^{n}\times\mathbb{R}^{l}}|\partial_{\rho}b(t,x,\rho,u)(x^{\prime},u^{\prime})|^{2}\rho(\mathrm{d}x^{\prime},\mathrm{d}u^{\prime})\right)^{\frac{1}{2}}$
$\qquad+\left(\int_{\mathbb{R}^{n}\times\mathbb{R}^{l}}|\partial_{\rho}\sigma(t,x,\rho,u)(x^{\prime},u^{\prime})|^{2}\rho(\mathrm{d}x^{\prime},\mathrm{d}u^{\prime})\right)^{\frac{1}{2}}\leq M,$ (21)
for the positive constant $M$ being given in Assumption 1-(A3).
Remark 3.4.
Compared with Acciaio et al. [1] or Hu et al. [26], we do not need to impose the uniform boundedness on gradients $\nabla_{u}(b,\sigma)$ of coefficients $(b,\sigma)$ of the state dynamics, which is relaxed to be the linear growth condition as in Assumption 1-(A4).
We then consider the following constrained extended MFC problem given by
$\begin{cases}\text{minimize $J(\alpha)$ over $\alpha\in\mathscr{U}[0,T]$;}\\[6.00006pt] \displaystyle\text{subject to $X_{t}^{\alpha}=\kappa+\int_{0}^{t}b(s,X_{s}^{\alpha},\mathcal{L}(X_{s}^{\alpha},\alpha_{s}),\alpha_{s})\mathrm{d}s$}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$+\int_{0}^{t}\sigma(s,X_{s}^{\alpha},\mathcal{L}(X_{s}^{\alpha},\alpha_{s}),\alpha_{s})\mathrm{d}W_{s}$,~{}~{}$\forall t\in[0,T]$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\mathbb{E}\left[\phi^{i}\left(t,X_{t}^{\alpha},\widetilde{\alpha}^{t}\right)\right]\geq 0$,~{}~{}$i\in I_{1}$,~{}~{}$\forall t\in[0,T]$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\psi^{j}\left(t,X_{t}^{\alpha},\alpha_{t}\right)\geq 0$,~{}~{}$j\in I_{2}$,~{}~{}$\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.}\end{cases}$ (22)
Obviously, it is equivalent to minimizing the objective functional $J(\alpha)$ in (22) over $\mathscr{U}_{ad}[0,T]$. We first have the following preliminary results on constraints whose proof is reported in Appendix A.
Lemma 3.5.
Let $i\in I_{1}$ and $(X,\alpha)\in\mathbb{M}^{2}\times\mathbb{L}_{\mathbb{F}}^{2}$. Define $g_{i}(t):=\mathbb{E}[\phi^{i}(t,X_{t},\widetilde{\alpha}^{t})]$ with $t\in[0,T]$. Then $g_{i}\in C([0,T];\mathbb{R})$.
Now, we introduce that
$\begin{cases}\displaystyle\mathtt{X}=\mathbb{M}^{2}\times\mathbb{L}_{\mathbb{F}}^{2},\\[5.0pt] \displaystyle\mathtt{Y}_{i}=C([0,T];\mathbb{R}),\quad\forall i\in I_{1},\\[5.0pt] \displaystyle\mathtt{Y}_{j}=\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}},\quad\forall j\in I_{2},\\[5.0pt] \displaystyle\mathtt{H}=\mathbb{M}^{2},\end{cases}$ (23)
where, for $x_{1}=(X_{1},\alpha_{1}),x_{2}=(X_{2},\alpha_{2})\in\mathtt{X}$, we define the inner product by
$\langle x_{1},x_{2}\rangle_{\mathtt{X}}=\langle X_{1},X_{2}\rangle_{\mathbb{M}^{2}}+\langle\alpha_{1},\alpha_{2}\rangle_{\mathbb{L}_{\mathbb{F}}^{2}}.$ (24)
Let $\lVert\cdot\rVert_{\mathtt{X}}:=\sqrt{\langle\cdot,\cdot\rangle_{\mathtt{X}}}$ be the induced norm. We also take $\|\cdot\|_{\mathtt{Y}_{i}}$ $(i\in I_{1})$, $\|\cdot\|_{\mathtt{Y}_{j}}$ $(j\in I_{2})$ and $\langle\cdot,\cdot\rangle_{\mathtt{H}}$ as the supremum norm $\|\cdot\|_{\infty}$ on $C([0,T];\mathbb{R}^{n})$, the norm $\|\cdot\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}$ (c.f. (15)) and the inner product $\langle\cdot,\cdot\rangle_{\mathbb{M}^{2}}$ (c.f. (14)), respectively. Furthermore, we define $J:\mathtt{X}\to\mathbb{R}$, $\Phi_{i}:\mathtt{X}\to\mathtt{Y}_{i}$  $(i\in I_{1})$, $\Psi_{j}:\mathtt{X}\to\mathtt{Y}_{j}$ $(j\in I_{2})$ and $F:\mathtt{X}\to\mathtt{H}$ as follows, for $(X,\alpha)\in\mathtt{X}$,
[…]
The mappings $\Phi_{i}$ ($i\in I_{1}$) and $\Psi_{j}~{}(j\in I_{2})$ in (25) are well-defined thanks to Assumption 1-(A2) and Lemma 3.5. Thus, we can formulate the following optimization problem with constraints that
$\begin{cases}\displaystyle\inf~{}J(X,\alpha)\\[3.99994pt] \displaystyle\mathrm{s.t.}~{}~{}(X,\alpha)\in C:=\mathbb{M}^{2}\times\mathscr{U}[0,T],\\[3.99994pt] \displaystyle~{}~{}~{}~{}~{}~{}\Phi_{i}(X,\alpha)\leq_{K_{1}}0,~{}~{}\forall i\in I_{1},\\[3.99994pt] \displaystyle~{}~{}~{}~{}~{}~{}\Psi_{j}(X,\alpha)\leq_{K_{2}}0,~{}~{}\forall j\in I_{2},\\[3.99994pt] \displaystyle~{}~{}~{}~{}~{}~{}F(X,\alpha)=0,\end{cases}$ (26)
Here $C=\mathbb{M}^{2}\times\mathscr{U}[0,T]$ is a closed and convex set in $\mathtt{X}$, and
$\begin{cases}\displaystyle K_{1}:=\{f\in C([0,T];\mathbb{R});~{}f(t)\geq 0,~{}\forall t\in[0,T]\},\\[6.00006pt] \displaystyle K_{2}:=\{f\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}};~{}f(t,\omega)\geq 0,~{}\forall(t,\omega)\in[0,T]\times\Omega\}=:\mathbb{L}_{\mathbb{F},\mathbb{R}+}^{2}\end{cases}$ (27)
are closed and convex cones in $C([0,T];\mathbb{R})$ and $\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}$, respectively. Moreover, one can verify that $K_{1}^{\rm o}=\{f\in C([0,T];\mathbb{R});~{}f(t)>0,~{}\forall t\in[0,T]\}$ and $K_{2}^{\rm o}=\varnothing$. On the other hand, the well-posedness of the controlled Mckean-Vlasov SDE (3.2) implies the following crucial equivalence whose proof is straightforward and hence omitted.
Lemma 3.6.
The value function of the primal control problem (22) and infimum of the new optimization problem (26) coincide. Moreover, if the minimizer exists for one problem, so does the other one and two minimizers also coincide.
Consequently, it is equivalent for us to consider the abstract optimization problem (26).
3.3 FJ optimality condition for abstract problem (26)
We next apply Theorem 2.7 to our C-(E)MFC problem in (26) and derive the FJ optimality condition. To this end, we need to provide the representations of Fréchet differentiablites of the mappings $J,~{}\Phi_{i}~{}(i\in I_{1}),~{}\Psi_{j}~{}(j\in I_{2})$ and $F$ defined in (25). In the sequel, let us set $\rho_{t}=\mathcal{L}(X_{t},\alpha_{t})$ for $t\in[0,T]$.
Lemma 3.7.
The mappings $J$, $\Phi_{i}$ ($i\in I_{1}$), $\Psi_{j}~{}(j\in I_{2})$ and $F$ defined in (25) are continuously Frêchet differentiable and their Frêchet derivatives have the following representations that, for $((X,\alpha),L,K)\in\mathtt{X}\times\mathbb{M}^{2}\times\mathbb{L}_{\mathbb{F}}^{2}$,
$D_{X}$ $J(X,\alpha)(L)=\mathbb{E}\left[\int_{0}^{T}\langle\nabla_{x}f(t,X_{t},\rho_{t},\alpha_{t}),L_{t}\rangle\mathrm{d}t+\langle\nabla_{x}g(X_{T},\mu_{T}),L_{T}\rangle\right.$
$\qquad+\left.\mathbb{E}^{\prime}\left[\int_{0}^{T}\langle\partial_{\mu}f(t,X_{t},\rho_{t},\alpha_{t})(X_{t}^{\prime},\alpha_{t}^{\prime}),L_{t}^{\prime}\rangle\mathrm{d}t+\langle\partial_{\mu}g(X_{T},\mu_{T})(X_{T}^{\prime}),L_{T}^{\prime}\rangle\right]\right],$
$D_{X}$ $\Phi_{i}(X,\alpha)(L)(t)=-\mathbb{E}[\langle\nabla_{x}\phi^{i}(t,X_{t},\widetilde{\alpha}^{t}),L_{t}\rangle],~{}~{}\forall t\in[0,T],~{}i\in I_{1},$
$D_{X}$ $\Psi_{i}(X,\alpha)(L)(t)=-\int_{0}^{T}\langle\nabla_{x}\psi^{j}(t,X_{t},\alpha_{t}),L_{t}\rangle~{}~{}\forall t\in[0,T],~{}j\in I_{2},$
$D_{X}$ $F(X,\alpha)(L)(\cdot)=L_{\cdot}-\int_{0}^{\cdot}\nabla_{x}b(s,X_{s},\rho_{s},\alpha_{s})L_{s}\mathrm{d}s-\int_{0}^{\cdot}\nabla_{x}\sigma(s,X_{s},\rho_{s},\alpha_{s})L_{s}\mathrm{d}W_{s}$
$\quad-\int_{0}^{\cdot}\mathbb{E}^{\prime}\left[\partial_{\mu}b(s,X_{s},\rho_{s},\alpha_{s})(X_{s}^{\prime},\alpha_{s}^{\prime})L_{s}^{\prime}\right]\mathrm{d}s-\int_{0}^{\cdot}\mathbb{E}^{\prime}\left[\partial_{\mu}\sigma(s,X_{s},\rho_{s},\alpha_{s})(X_{s}^{\prime},\alpha_{s}^{\prime})L_{s}^{\prime}\right]\mathrm{d}W_{s},$
$D_{\alpha}$ $J(X,\alpha)(K)=\mathbb{E}\left[\int_{0}^{T}\langle\nabla_{u}f(t,X_{t},\rho_{t},\alpha_{t}),K_{t}\rangle\mathrm{d}t+\int_{0}^{T}\langle\partial_{\nu}f(t,X_{t},\rho_{t},\alpha_{t})(X_{t}^{\prime},\alpha_{t}^{\prime}),K_{t}^{\prime}\rangle\mathrm{d}t\right],$
$D_{\alpha}$ $\Phi_{i}(X,\alpha)(K)(t)=-\mathbb{E}[\langle\nabla_{\beta}\phi^{i}(t,X_{t},\widetilde{\alpha}^{t}),\widetilde{K}^{t}\rangle_{\mathcal{X}}],~{}~{}\forall t\in[0,T],~{}i\in I_{1},$
$D_{\alpha}$ $\Psi_{i}(X,\alpha)(K)(t)=-\langle\nabla_{u}\psi^{j}(t,X_{t},\alpha_{t}),K_{t}\rangle,~{}~{}\forall t\in[0,T],~{}j\in I_{2},$
$D_{\alpha}$ $F(X,\alpha)(K)(\cdot)=-\int_{0}^{\cdot}\nabla_{u}b(s,X_{s},\rho_{s},\alpha_{s})K_{s}\mathrm{d}s-\int_{0}^{\cdot}\nabla_{u}\sigma(s,X_{s},\rho_{s},\alpha_{s})K_{s}\mathrm{d}W_{s}$
$\quad-\int_{0}^{\cdot}\mathbb{E}^{\prime}\left[\partial_{\nu}b(s,X_{s},\rho_{s},\alpha_{s})(X_{s}^{\prime},\alpha_{s}^{\prime})K_{s}^{\prime}\right]\mathrm{d}s-\int_{0}^{\cdot}\mathbb{E}^{\prime}\left[\partial_{\nu}\sigma(s,X_{s},\rho_{s},\alpha_{s})(X_{s}^{\prime},\alpha_{s}^{\prime})K_{s}^{\prime}\right]\mathrm{d}W_{s}.$
Here, we have used the notation $Mz:=(M_{1}z,\ldots,M_{r}z)\in\mathbb{R}^{n\times r}$ for any $M=(M_{1},\ldots,M_{r})\in\mathbb{R}^{n\times n\times r}$ and $z\in\mathbb{R}^{n}$, and $(X^{\prime},\alpha^{\prime},L^{\prime},K^{\prime})$ is an independent copy of $(X,\alpha,L,K)$ defined on the probability space $(\Omega^{\prime},\mathcal{F}^{\prime},P^{\prime})$ constructed following Lemma 3.2.
Proof.
The result is a direct consequence from simple calculations together with Lemma 3.2 and the fact that $\|L\|_{\mathbb{L}_{\mathbb{F}}^{2}}\leq\|L\|_{\mathbb{M}^{2}}$. ∎
We next provide the well-posed lemma whose proof is delegated into Appendix A.
Lemma 3.8.
Let $(B,\Sigma):[0,T]\times\Omega\times\mathbb{R}^{n}\times L^{2}(\Omega;\mathbb{R}^{n})\mapsto\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ be progressively measurable mappings and satisfy that, there exists a constant $M>0$ such that, for all $(t,\omega)\in[0,T]\times\Omega$ and $(x,X),(y,Y)\in\mathbb{R}^{n}\times L^{2}(\Omega;\mathbb{R}^{n})$,
$|(B,\Sigma)(t,\omega,x,X)-(B,\Sigma)(t,\omega,y,Y)|\leq M\left\{|x-y|+\mathbb{E}[|X-Y|^{2}]\right\},$ (28)
and for any $R>0$,
$\mathbb{E}\left[\int_{0}^{T}\sup_{|x|^{2}+\mathbb{E}[|X|^{2}]\leq R}\left(|B(t,\cdot,x,X)|^{2}+|\Sigma(t,\cdot,x,X)|^{2}\right)\mathrm{d}t\right]<+\infty.$ (29)
Then, the following (self-dependent) SDE admits a unique strong solution, for $t\in[0,T]$,
$\mathrm{d}Y_{t}=B(t,\omega,Y_{t},\overline{Y}_{t})\mathrm{d}t+\Sigma(t,\omega,Y_{t},\overline{Y}_{t})\mathrm{d}W_{t},\quad Y_{0}\in L^{2}((\Omega,\mathcal{F}_{0},\mathbb{P});\mathbb{R}^{n}).$ (30)
Here, for $t\in[0,T]$, $\overline{Y}_{t}$ stands for the random variable $Y_{t}$ instead of the evaluation $Y_{t}(\omega)$ at $\omega\in\Omega$.
It is not difficult to see that once the solution $Y=(Y_{t})_{t\in[0,T]}$ to the SDE (30) exists, it automatically belongs to $\mathbb{M}^{2}$ due to the square-integrable condition (29). Then, in lieu of Lemma 3.8, we have the next result.
Lemma 3.9.
For any $(X,\alpha)\in\mathtt{X}$, the mapping $D_{X}F(X,\alpha):\mathbb{M}^{2}\mapsto\mathbb{M}^{2}$ is bijective.
Proof.
To prove the surjection, it suffices to show that, for any $\zeta\in\mathbb{M}^{2}$ admitting the representation $\zeta_{t}=\zeta_{0}+\int_{0}^{t}\zeta^{1}_{s}\mathrm{d}s+\int_{0}^{t}\zeta^{2}_{s}\mathrm{d}W_{s}$ for $t\in[0,T]$, there exists a strong solution $L=(L_{t})_{t\in[0,T]}$ to the SDE, for $t\in[0,T]$,
$L_{t}$ $=\zeta_{0}+\int_{0}^{t}(\nabla_{x}b(s,X_{s},\rho_{s},\alpha_{s})L_{s}+\zeta_{s}^{1})\mathrm{d}s+\int_{0}^{t}(\nabla_{x}\sigma(s,X_{s},\rho_{s},\alpha_{s})L_{s}+\zeta_{s}^{2})\mathrm{d}W_{s}$ (31)
$\quad+\int_{0}^{t}\mathbb{E}^{\prime}\left[\partial_{\mu}b(s,X_{s},\rho_{s},\alpha_{s})(X_{s}^{\prime},\alpha_{s}^{\prime})L_{s}^{\prime}\right]\mathrm{d}s+\int_{0}^{t}\mathbb{E}^{\prime}\left[\partial_{\mu}\sigma(s,X_{s},\rho_{s},\alpha_{s})(X_{s}^{\prime},\alpha_{s}^{\prime})L_{s}^{\prime}\right]\mathrm{d}W_{s}.$
In fact, the existence holds if we define $(B,\Sigma)$ in Lemma 3.8 by, for $(t,\omega,x,X)\in[0,T]\times\Omega\times\mathbb{R}^{n}\times L^{2}(\Omega;\mathbb{R}^{n})$,
$\begin{cases}\displaystyle B(t,\omega,x,X):=\nabla_{x}b(t,X_{t}(\omega),\rho_{t},\alpha_{t}(\omega))x+\mathbb{E}^{\prime}\left[\partial_{\mu}b(t,X_{t}(\omega),\rho_{t},\alpha_{t}(\omega))(X_{s}^{\prime},\alpha_{s}^{\prime})X^{\prime}\right]+\zeta_{t}^{1}(\omega),\\[8.00003pt] \Sigma(t,\omega,x,X):=\nabla_{x}\sigma(t,X_{t}(\omega),\rho_{t},\alpha_{t}(\omega))x+\mathbb{E}^{\prime}\left[\partial_{\mu}\sigma(t,X_{t}(\omega),\rho_{t},\alpha_{t}(\omega))(X_{s}^{\prime},\alpha_{s}^{\prime})X^{\prime}\right]+\zeta_{t}^{2}(\omega),\end{cases}$
and by exploiting (3.3), it holds that
$\sup_{(t,x)\in[0,T]\times\mathbb{R}^{n}}\left(\nabla_{x}b(t,X_{t},\rho_{t},\alpha_{t})|+|\nabla_{x}(t,X_{t},\rho_{t},\alpha_{t})|\right)\leq M,~{}{\rm and}$
$\qquad\qquad\qquad\mathbb{E}\left[|\zeta_{0}|^{2}+\int_{0}^{T}\left(|\zeta_{t}^{1}|^{2}+|\zeta_{t}^{2}|^{2}\right)\mathrm{d}t\right]<\infty.$
On the other hand, the injection result holds given that $\ker(D_{X}F(X,\alpha))=0$, which is equivalent to show that the SDE, for $t\in[0,T]$,
$L_{t}=\int_{0}^{t}\nabla_{x}b(s,X_{s},\rho_{s},\alpha_{s})L_{s}\mathrm{d}s+\int_{0}^{t}\nabla_{x}\sigma(s,X_{s},\rho_{s},\alpha_{s})L_{s}\mathrm{d}W_{s}$
admits only a zero solution, which is also a direct consequence of Lemma 3.8. Hence, the proof of the lemma is complete. ∎
Recall the closed and convex cones $K_{1}$ and $K_{2}$ introduced in (27). It follows from Riesz representation theorem that
$K_{1}^{+}=\mathcal{M}^{+}[0,T],\quad K_{2}^{+}=\{f\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}};~{}f(t,\omega)\geq 0,~{}\forall(t,\omega)\in[0,T]\times\Omega\}=\mathbb{L}_{\mathbb{F},\mathbb{R}+}^{2}.$ (32)
The next result follows from Theorem 2.7 together with (23) and (32).
Lemma 3.10 (FJ optimality condition for problem (26)).
Let $(\widehat{X},\widehat{\alpha})\in C=\mathbb{M}^{2}\times\mathscr{U}[0,T]$ be an optimal pair to the C-(E)MFC problem (22) (equivalent to the abstract formulation (26)). Then, there exist (nontrivial) multipliers $r_{0}\geq 0$, $\mu^{i}\in\mathcal{M}^{+}[0,T]$ for $i\in I_{1}$, $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $j\in I_{2}$ and $\lambda\in\mathbb{M}^{2}$ which are not all zeros such that
$\begin{cases}\displaystyle\int_{[0,T]}\Phi_{i}(\widehat{X},\widehat{\alpha})(t)\mu^{i}(\mathrm{d}t)=0,~{}~{}\forall i\in I_{1},\\[13.99995pt] \displaystyle\mathbb{E}\left[\int_{0}^{T}\Psi_{j}(\widehat{X},\widehat{\alpha})(t)\eta^{j}_{t}\mathrm{d}t\right]=0,~{}~{}\forall j\in I_{2},\\[13.99995pt] \displaystyle r_{0}+\sum_{i\in I_{1}}\|\mu^{i}\|_{\rm TV}+\sum_{j\in I_{2}}\|\eta^{j}\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}+\|\lambda\|_{\mathbb{M}^{2}}=1,\\[13.99995pt] \displaystyle-\xi_{1}=r_{0}D_{X}J(\widehat{X},\widehat{\alpha})(\cdot)+\sum_{i\in I_{1}}\int_{[0,T]}D_{X}\Phi_{i}(\widehat{X},\widehat{\alpha})(\cdot)(t)\mu^{i}(\mathrm{d}t)\\[8.00003pt] \displaystyle\quad\qquad+\sum_{j\in I_{2}}\mathbb{E}\left[\int_{0}^{T}D_{X}\Psi_{j}(\widehat{X},\widehat{\alpha})(\cdot)(t)\eta^{j}_{t}\mathrm{d}t\right]-\left\langle\lambda,D_{X}F(\widehat{X},\widehat{\alpha})(\cdot)\right\rangle_{\mathbb{M}^{2}}~{}~{}\text{\rm in}~{}\mathbb{M}^{2},\\[13.99995pt] \displaystyle-\xi_{2}=r_{0}D_{\alpha}J(\widehat{X},\widehat{\alpha})(\cdot)+\sum_{i\in I_{1}}\int_{[0,T]}D_{\alpha}\Phi_{i}(\widehat{X},\widehat{\alpha})(\cdot)(t)\mu^{i}(\mathrm{d}t)\\[8.00003pt] \displaystyle\quad\qquad+\sum_{j\in I_{2}}\mathbb{E}\left[\int_{0}^{T}D_{\alpha}\Psi_{j}(\widehat{X},\widehat{\alpha})(\cdot)(t)\eta^{j}_{t}\mathrm{d}t\right]-\left\langle\lambda,D_{\alpha}F(\widehat{X},\widehat{\alpha})(\cdot)\right\rangle_{\mathbb{M}^{2}}~{}~{}\text{\rm in}~{}\mathbb{L}_{\mathbb{F}}^{2},\end{cases}$ (33)
where $\|\cdot\|_{\rm TV}$ denotes the total variation norm on the space of all (signed) Radon measures and the element $\xi=(\xi_{1},\xi_{2})\in N_{C}(\widehat{X},\widehat{\alpha})\subset\mathtt{X}=\mathbb{M}^{2}\times\mathbb{L}_{\mathbb{F}}^{2}$.
Building upon the FJ optimality condition for the abstract problem (26), we obtain the stochastic maximum principle (SMP) for our C-(E)MFC problems, which is another main result of this paper.
Theorem 3.11 (SMP for extended MFC problem with constraints).
Let $(\widehat{X},\widehat{\alpha})\in C:=\mathbb{M}^{2}\times\mathscr{U}[0,T]$ be an optimal solution to our C-(E)MFC problem (22). There exist some $r_{0}\geq 0$, $\mu^{i}\in\mathcal{M}^{+}[0,T]$ for $i\in I_{1}$ and $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $j\in I_{2}$ such that
  • (i)
    the Radon measure $\mu^{i}$ is supported on $\{t\in[0,T];~{}\mathbb{E}[\phi^{i}(t,\widehat{X}_{t},\widetilde{\widehat{\alpha}}^{t})]=0\}$ for all $i\in I_{1}$,
  • (ii)
    for $j\in I_{2}$, since $\eta_{t}^{j}\geq 0$, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-$\mathrm{a.s.}$ and $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$, we can define the (continuous) increasing process $A^{j}=(A_{t}^{j})_{t\in[0,T]}$ by $A_{t}^{j}=\int_{0}^{t}\eta_{s}^{j}\mathrm{d}s$. Then, for $\mathbb{P}$-$\mathrm{a.s.}$ $\omega$, the process $A^{j}(\omega)$ only increases on the set $\{t\in[0,T];~{}\psi^{j}(t,\widehat{X}_{t}(\omega),\widehat{\alpha}_{t}(\omega))=0\}$, $\mathrm{d}t$-$\mathrm{a.s.}$,
  • (iii)
    it holds that $r_{0}+\sum_{i\in I_{1}}\mu^{i}([0,T])+\sum_{j\in I_{2}}\|\eta^{j}\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}=1$.
Furthermore, the stochastic (first-order) minimum condition holds that, for any $u\in U$ and $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-$\mathrm{a.s.}$,
$\left\langle\nabla_{u}H^{r_{0}}(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t},Y_{t},Z_{t})+\mathbb{E}^{\prime}\left[\partial_{\nu}H^{r_{0}}(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime},Y_{t},Z_{t})(\widehat{X}_{t},\widehat{\alpha}_{t})\right]\right.$
$\qquad\left.-\sum_{i\in I_{1}}\int_{t}^{T}D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}\mu^{i}(\mathrm{d}s)-\sum_{j\in I_{2}}\nabla_{u}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta_{t}^{j},u-\widehat{\alpha}_{t}\right\rangle\geq 0.$ (34)
Here, $D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}$ is the evaluation of the continuous function $D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})$ at time $t$ and $\widehat{\rho}_{t}:={\cal L}(\widehat{X}_{t},\widehat{\alpha}_{t})$ for $t\in[0,T]$. The Hamiltonian $H^{r_{0}}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathbb{R}^{l}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\mapsto\mathbb{R}$ is defined by
$H^{r_{0}}(t,x,\rho,u,y,z):=\langle b(t,x,\rho,u),y\rangle+\mathrm{tr}\left(\sigma(t,x,\rho,u)z^{\top}\right)+r_{0}f(t,x,\rho,u),$ (35)
and the process pair $(Y,Z)=(Y_{t},Z_{t})_{t\in[0,T]}$ taking values in $\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ is the unique solution to the following general type of constrained BSDE:
$Y_{t}$ $=r_{0}\nabla_{x}g(\widehat{X}_{T},\widehat{\mu}_{T})+r_{0}\mathbb{E}^{\prime}\left[\partial_{\mu}g(\widehat{X}_{T},\widehat{\mu}_{T})(\widehat{X}_{T}^{\prime})\right]+\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\rho}_{s},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s$
$\quad+\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}H^{r_{0}}(s,\widehat{X}_{s}^{\prime},\widehat{\rho}_{s},\widehat{\alpha}_{s}^{\prime},Y_{s},Z_{s})(\widehat{X}_{s},\widehat{\alpha}_{s})\right]\mathrm{d}s-\sum_{i\in I_{1}}\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)$
$\quad-\sum_{j\in I_{2}}\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\alpha}_{s})\mathrm{d}A_{s}^{j}-\int_{t}^{T}Z_{s}\mathrm{d}W_{s}$ (36)
with $\widehat{\mu}_{T}=\mathcal{L}(\widehat{X}_{T})$. Here, $(\widehat{X}^{\prime},\widehat{\alpha}^{\prime})$ is an independent copy of $(\widehat{X},\widehat{\alpha})$ defined on the probability space $(\Omega^{\prime},\mathcal{F}^{\prime},P^{\prime})$ constructed following Lemma 3.2.
Remark 3.12.
In terms of (35), when $r_{0}=1$, our Hamiltonian $H^{r_{0}}$ reduces to the classical Hamiltonian. In particular, when $m=0$ (i.e., there are no inequality constraints in problem (22)), we have $r_{0}=1$ by using Corollary 2.13, Lemma 3.9 (or simply Theorem 3.11-(iii)), and hence Theorem 3.11 reduces to the classical stochastic maximum principle of the extended MFC problem (c.f. Theorem 3.2 of Acciaio et al. [1]).
Proof of Theorem 3.11.
By using the definition (3) of normal cone $N_{C}(\widehat{X},\widehat{\alpha})$, it follows that, for all $(X,\alpha)\in C:=\mathbb{M}^{2}\times\mathscr{U}[0,T]$ and $\xi=(\xi_{1},\xi_{2})\in N_{C}(\widehat{X},\widehat{\alpha})$,
$\langle\xi_{1},\widehat{X}-X\rangle_{\mathbb{M}^{2}}+\langle\xi_{2},\widehat{\alpha}-\alpha\rangle_{\mathbb{L}_{\mathbb{F}}^{2}}\geq 0.$
Taking $\alpha=\widehat{\alpha}$. Then, we can conclude that $\xi_{1}=0$, and hence $\xi_{2}\in N_{\mathscr{U}[0,T]}(\widehat{\alpha})$, i.e.,
$\langle\xi_{2},\widehat{\alpha}-\alpha\rangle_{\mathbb{L}_{\mathbb{F}}^{2}}\geq 0,~{}~{}\forall\alpha\in\mathscr{U}[0,T].$ (37)
We can thus rewrite the fourth equality documented in (33) of Lemma 3.10 as follows:
00 0 $=r_{0}D_{X}J(\widehat{X},\widehat{\alpha})(\cdot)+\sum_{i\in I_{1}}\int_{[0,T]}D_{X}\Phi_{i}(\widehat{X},\widehat{\alpha})(\cdot)(t)\mu^{i}(\mathrm{d}t)$
$\quad+\sum_{j\in I_{2}}\mathbb{E}\left[\int_{0}^{T}D_{X}\Psi_{j}(\widehat{X},\widehat{\alpha})(\cdot)(t)\eta^{j}_{t}\mathrm{d}t\right]-\left\langle\lambda,D_{X}F(\widehat{X},\widehat{\alpha})(\cdot)\right\rangle_{\mathbb{M}^{2}}.$ (38)
We claim that $r_{0}$, $\mu^{i}$ for $i\in I_{1}$ and $\eta^{j}$ for $j\in I_{2}$ can not be zeros simultaneously, since otherwise, the equality (3.3) yields that, for all $L\in\mathbb{M}^{2}$ with the nonzero multiplier $\lambda\in\mathbb{M}^{2}$,
$\left\langle\lambda,D_{X}F(\widehat{X},\widehat{\alpha})(L)\right\rangle_{\mathbb{M}^{2}}=0.$
Then, it follows from Lemma 3.9 that $\lambda=0$, which creates a contradiction.
Now, we can make the scaling on both sides of the equality (3.3) by multiplying a positive constant $r>0$ such that $r(|r_{0}|+\sum_{i\in I_{1}}\|\mu^{i}\|_{\rm TV}+\sum_{j\in I_{2}}\|\eta^{j}\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}})=1$. To ease the notation, we still denote by $r_{0}$, $\mu^{i}$ for $i\in I_{1}$, $\eta^{j}$ for $j\in I_{2}$ and $\lambda$ the scaled multipliers respectively. That is to say
$|r_{0}|+\sum_{i\in I_{1}}\|\mu^{i}\|_{\rm TV}+\sum_{j\in I_{2}}\|\eta^{j}\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}=1.$
For any $L\in\mathbb{M}^{2}$, by inserting the corresponding Frêchet derivatives given in Lemma 3.7 into (3.3), and let the Itô’s representation of $(\lambda,L)$ be given by $(\lambda_{t},L_{t})=(\lambda_{0},L_{0})+\int_{0}^{t}(\lambda_{s}^{1},L_{s}^{1})\mathrm{d}s+\int_{0}^{t}(\lambda_{s}^{2},L_{s}^{2})\mathrm{d}W_{s}$ with $t\in[0,T]$ since $\lambda,L\in\mathbb{M}^{2}$, we arrive at
00 0 $=r_{0}\mathbb{E}\left[\int_{0}^{T}\langle\nabla_{x}f(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t}),L_{t}\rangle\mathrm{d}t+\langle\nabla_{x}g(\widehat{X}_{T},\widehat{\mu}_{T}),L_{T}\rangle\right.$
$\quad+\left.\mathbb{E}^{\prime}\left[\int_{0}^{T}\langle\partial_{\mu}f(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime},\widehat{\alpha}_{t}^{\prime}),L_{t}^{\prime}\rangle\mathrm{d}t+\langle\partial_{\mu}g(\widehat{X}_{T},\widehat{\mu}_{T})(\widehat{X}_{T}^{\prime}),L_{T}^{\prime}\rangle\right]\right]$
$\quad-\sum_{i\in I_{1}}\int_{[0,T]}\mathbb{E}[\langle\nabla_{x}\phi^{i}(t,\widehat{X}_{t},\widetilde{\widehat{\alpha}}^{t}),L_{t}\rangle]\mu^{i}(\mathrm{d}t)-\sum_{j\in I_{2}}\mathbb{E}\left[\int_{0}^{T}\langle\nabla_{x}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t}),L_{t}\rangle\eta^{j}_{t}\mathrm{d}t\right]$ (39)
$\quad-\mathbb{E}\left[\langle\lambda_{0},L_{0}\rangle+\int_{0}^{T}\langle\lambda_{t}^{1},L_{t}^{1}-\nabla_{x}b(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})L_{t}\rangle\mathrm{d}t+\int_{0}^{T}\mathrm{tr}\left(\lambda_{t}^{2}(L_{t}^{2}-\nabla_{x}\sigma(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})L_{t})^{\top}\right)\mathrm{d}t\right.$
$\quad+\left.\int_{0}^{T}\left\langle\lambda_{t}^{1},\mathbb{E}^{\prime}\left[\partial_{\mu}b(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime},\widehat{\alpha}_{t}^{\prime})L_{t}^{\prime}\right]\right\rangle\mathrm{d}t+\int_{0}^{T}\mathrm{tr}\left(\lambda_{t}^{2}\left(\mathbb{E}^{\prime}\left[\partial_{\mu}\sigma(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime},\widehat{\alpha}_{t}^{\prime})L_{t}^{\prime}\right]\right)^{\top}\right)\mathrm{d}t\right]$
with $\widehat{\rho}_{t}=\mathcal{L}(\widehat{X}_{t},\widehat{\alpha}_{t})$ for $t\in[0,T]$ and $\widehat{\mu}_{T}=\mathcal{L}(\widehat{X}_{T})$. We then define an element $p=(p_{t})_{t\in[0,T]}$ in $\mathbb{M}^{2}$ as follows, for $t\in[0,T]$,
$p_{t}$ $=r_{0}\nabla_{x}g(\widehat{X}_{T},\widehat{\mu}_{T})+r_{0}\mathbb{E}^{\prime}\left[\partial_{\mu}g(\widehat{X}_{T}^{\prime},\widehat{\mu}_{T})(\widehat{X}_{T})\right]$
$\quad-\sum_{i\in I_{1}}\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)-\sum_{j\in I_{2}}\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\alpha}_{s})\eta^{j}_{s}\mathrm{d}s$
$\quad+\int_{t}^{T}\left[r_{0}\nabla_{x}f(s,\widehat{X}_{s},\widehat{\rho}_{s},\widehat{\alpha}_{s})+\nabla_{x}b(s,\widehat{X}_{s},\widehat{\rho}_{s},\widehat{\alpha}_{s})\lambda_{s}^{1}+\nabla_{x}\sigma(s,\widehat{X}_{s},\widehat{\rho}_{s},\widehat{\alpha}_{s})\lambda_{s}^{2}\right]\mathrm{d}s$
$\quad+\int_{t}^{T}\mathbb{E}^{\prime}\left[r_{0}\partial_{\mu}f(s,\widehat{X}_{s}^{\prime},\widehat{\rho}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s},\widehat{\alpha}_{s})+\partial_{\mu}b(s,\widehat{X}_{s}^{\prime},\widehat{\rho}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s},\widehat{\alpha}_{s}){\lambda_{s}^{1}}^{\prime}\right.$
$\qquad\qquad+\left.\partial_{\mu}\sigma(s,\widehat{X}_{s}^{\prime},\widehat{\rho}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s},\widehat{\alpha}_{s}){\lambda_{s}^{2}}^{\prime}\right]\mathrm{d}s-\int_{t}^{T}\lambda_{s}^{2}\mathrm{d}W_{s},$ (40)
where $MN:=\sum_{i=1}^{r}M_{i}L_{i}\in\mathbb{R}^{n}$ for any $M=(M_{1},\ldots,M_{r})\in\mathbb{R}^{n\times n\times r}$ and $L=(L_{1},\ldots,L_{r})\in\mathbb{R}^{n\times r}$. Here, the integral $\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)$ is well-defined because $\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})$ is Borel measurable w.r.t. time $s$ and $\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)$ is finite $\mathbb{P}$-$\mathrm{a.s.}$, which results from $\mathbb{E}[\int_{[t,T]}|\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})|\mu^{i}(\mathrm{d}s)]<\infty$ for $i\in I_{1}$. In addition, $\mathbb{E}[\int_{0}^{T}\nabla_{x}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta_{t}^{j}\mathrm{d}t]<\infty$, and hence $\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\alpha}_{s})\eta_{s}^{j}\mathrm{d}s$ is $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-$\mathrm{a.s.}$ well-defined. Applying Itô’s formula, we can derive that
$\mathbb{E}\left[\int_{0}^{T}\langle\lambda_{t}^{1},L_{t}^{1}\rangle\mathrm{d}t\right]=\mathbb{E}\left[\int_{0}^{T}\langle p_{t},\mathrm{d}L_{t}\rangle+\int_{0}^{T}\langle\lambda_{t}^{1}-p_{t},\mathrm{d}L_{t}\rangle\right]$
$\quad=\mathbb{E}\left[\int_{0}^{T}\langle\lambda_{t}^{1}-p_{t},\mathrm{d}L_{t}\rangle+\langle p_{T},L_{T}\rangle-\langle p_{0},L_{0}\rangle-\int_{0}^{T}\langle\mathrm{d}p_{t},L_{t}\rangle-\langle p,L\rangle_{T}\right]$
$=\mathbb{E}\left[\int_{0}^{T}\langle\lambda_{t}^{1}-p_{t},\mathrm{d}L_{t}\rangle\right]+\mathbb{E}\left[\left\langle r_{0}\nabla_{x}g(\widehat{X}_{T},\widehat{\mu}_{T}),L_{T}\right\rangle+\left\langle r_{0}\mathbb{E}^{\prime}\left[\partial_{\mu}g(\widehat{X}_{T}^{\prime},\widehat{\mu}_{T})(\widehat{X}_{T})\right],L_{T}\right\rangle\right]$
$\quad+\mathbb{E}\left[\int_{0}^{T}\left\langle r_{0}\nabla_{x}f(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t}),L_{t}\right\rangle\mathrm{d}t+\int_{0}^{T}\left\langle r_{0}\mathbb{E}^{\prime}\left[\partial_{\mu}f(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})\right],L_{t}\right\rangle\mathrm{d}t\right]$
$\quad+\mathbb{E}\left[\int_{0}^{T}\left(\langle\lambda_{t}^{1},\nabla_{x}b(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})L_{t}\rangle+\mathrm{tr}\left(\lambda_{t}^{2}\left(\nabla_{x}\sigma(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})L_{t}-L_{t}^{2}\right)^{\top}\right)\right)\right]$
$\quad+\mathbb{E}\left[\int_{0}^{T}\left\langle\lambda_{t}^{1},\mathbb{E}^{\prime}\left[\partial_{\mu}b(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})L_{t}\right]\right\rangle\mathrm{d}t-\langle p_{0},L_{0}\rangle\right].$
$\quad+\mathbb{E}\left[\int_{0}^{T}\mathrm{tr}\left(\lambda_{t}^{2}\left(\mathbb{E}^{\prime}\left[\partial_{\mu}\sigma(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})L_{t}\right]\right)^{\top}\right)\mathrm{d}t\right]$
$\quad-\sum_{i\in I_{1}}\int_{[0,T]}\mathbb{E}[\langle\nabla_{x}\phi^{i}(t,\widehat{X}_{t},\widetilde{\widehat{\alpha}}^{t}),L_{t}\rangle]\mu^{i}(\mathrm{d}t)-\sum_{j\in I_{2}}\int_{0}^{T}\mathbb{E}\left[\langle\nabla_{x}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta_{t}^{j},L_{t}\rangle\right]\mathrm{d}t$
$=\mathbb{E}\left[\int_{0}^{T}\langle\lambda_{t}^{1}-p_{t},\mathrm{d}L_{t}\rangle\right]+\mathbb{E}\left[\left\langle r_{0}\nabla_{x}g(\widehat{X}_{T},\widehat{\mu}_{T}),L_{T}\right\rangle+\mathbb{E}^{\prime}\left[\left\langle r_{0}\partial_{\mu}g(\widehat{X}_{T},\widehat{\mu}_{T})(\widehat{X}_{T}^{\prime}),L_{T}^{\prime}\right\rangle\right]\right]$
$\quad+\mathbb{E}\left[\int_{0}^{T}\left\langle r_{0}\nabla_{x}f(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t}),L_{t}\right\rangle\mathrm{d}t+\int_{0}^{T}\left\langle r_{0}\mathbb{E}^{\prime}\left[\partial_{\mu}f(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})\right],L_{t}\right\rangle\mathrm{d}t\right]$
$\quad+\mathbb{E}\left[\int_{0}^{T}\left(\langle\lambda_{t}^{1},\nabla_{x}b(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})L_{t}\rangle+\mathrm{tr}\left(\lambda_{t}^{2}\left(\nabla_{x}\sigma(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})L_{t}-L_{t}^{2}\right)^{\top}\right)\right)\right]$
$\quad+\mathbb{E}\left[\int_{0}^{T}\left\langle\lambda_{t}^{1},\mathbb{E}^{\prime}\left[\partial_{\mu}b(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime},\widehat{\alpha}_{t}^{\prime})L_{t}^{\prime}\right]\right\rangle\mathrm{d}t-\langle p_{0},L_{0}\rangle\right].$
$\quad+\mathbb{E}\left[\int_{0}^{T}\mathrm{tr}\left(\lambda_{t}^{2}\left(\mathbb{E}^{\prime}\left[\partial_{\mu}\sigma(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime},\widehat{\alpha}_{t}^{\prime})L_{t}^{\prime}\right]\right)^{\top}\right)\mathrm{d}t\right]$
$\quad-\sum_{i\in I_{1}}\int_{[0,T]}\mathbb{E}[\langle\nabla_{x}\phi^{i}(t,\widehat{X}_{t},\widetilde{\widehat{\alpha}}^{t}),L_{t}\rangle]\mu^{i}(\mathrm{d}t)-\sum_{j\in I_{2}}\int_{0}^{T}\mathbb{E}\left[\langle\nabla_{x}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta_{t}^{j},L_{t}\rangle\right]\mathrm{d}t,$
where $\langle p,L\rangle_{T}$ stands for the quadratic variation of $p$ and $L$, and we have applied the Fubini theorem. Plugging the last equality back into (3.3), we can conclude that
$\mathbb{E}\left[\langle p_{0}-\lambda_{0},L_{0}\rangle+\int_{0}^{T}\langle p_{t}-\lambda_{t}^{1},\mathrm{d}L_{t}\rangle\right]=0,~{}~{}\forall L\in\mathbb{M}^{2}.$
By the arbitrariness of the element $L\in\mathbb{M}^{2}$ and the continuity of $t\mapsto p_{t}$, we can derive that $p_{0}=\lambda_{0}$, $\mathbb{P}$-$\mathrm{a.s.}$, and $p$ is indistinguishable from $\lambda^{1}=(\lambda_{t}^{1})_{t\in[0,T]}$. Together with (3.3), this results in that, for $t\in[0,T]$,
$\lambda_{t}^{1}$ $=r_{0}\nabla_{x}g(\widehat{X}_{T},\widehat{\mu}_{T})+r_{0}\mathbb{E}^{\prime}\left[\partial_{\mu}g(\widehat{X}_{T}^{\prime},\widehat{\mu}_{T})(\widehat{X}_{T})\right]$
$\quad-\sum_{i\in I_{1}}\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)-\sum_{j\in I_{2}}\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\alpha}_{s})\eta^{j}_{s}\mathrm{d}s$
$\quad+\int_{t}^{T}\left[r_{0}\nabla_{x}f(s,\widehat{X}_{s},\widehat{\rho}_{s},\widehat{\alpha}_{s})+\nabla_{x}b(s,\widehat{X}_{s},\widehat{\rho}_{s},\widehat{\alpha}_{s})\lambda_{s}^{1}+\nabla_{x}\sigma(s,\widehat{X}_{s},\widehat{\rho}_{s},\widehat{\alpha}_{s})\lambda_{s}^{2}\right]\mathrm{d}s$
$\quad+\int_{t}^{T}\mathbb{E}^{\prime}\left[r_{0}\partial_{\mu}f(s,\widehat{X}_{s}^{\prime},\widehat{\rho}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s},\widehat{\alpha}_{s})+\partial_{\mu}b(s,\widehat{X}_{s}^{\prime},\widehat{\rho}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s},\widehat{\alpha}_{s}){\lambda_{s}^{1}}^{\prime}\right.$
$\qquad\qquad\quad+\left.\partial_{\mu}\sigma(s,\widehat{X}_{s}^{\prime},\widehat{\rho}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s},\widehat{\alpha}_{s}){\lambda_{s}^{2}}^{\prime}\right]\mathrm{d}s-\int_{t}^{T}\lambda_{s}^{2}\mathrm{d}W_{s}.$ (41)
Here, we recall that $(\widehat{X}^{\prime},\widehat{\alpha}^{\prime},\lambda^{1^{\prime}},\lambda^{2^{\prime}})$ is an independent copy of $(\widehat{X},\widehat{\alpha},\lambda^{1},\lambda^{2})$ defined on the probability space $(\Omega^{\prime},\mathcal{F}^{\prime},P^{\prime})$ constructed following Lemma 3.2.
Now, if we set $Q:\mathbb{M}^{2}\mapsto\mathbb{M}^{2}$ that maps any $\eta\in\mathbb{M}^{2}$ admitting the Itô’s representation $\eta_{t}=\eta_{0}+\int_{0}^{t}\eta_{s}^{1}\mathrm{d}s+\int_{0}^{t}\eta_{s}^{2}\mathrm{d}W_{s}$ with $t\in[0,T]$ to $Q(\eta)(\cdot):=\eta_{0}+\int_{0}^{\cdot}\eta_{s}\mathrm{d}s+\int_{0}^{\cdot}\eta_{s}^{2}\mathrm{d}W_{s}$ for $t\in[0,T]$. Then, we can deduce that $Q(\lambda^{1})$ is indistinguishable from $\lambda$. Moreover, combining the Frêchet derivative provided in Lemma 3.7 together with (37) and the fifth equality in (33) on the FJ condition in Lemma 3.10, we get that, for any $\alpha\in\mathscr{U}[0,T]$,
00 0 $\leq\mathbb{E}\left[\int_{0}^{T}r_{0}\left\langle\nabla_{u}f(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})+\mathbb{E}^{\prime}\left[\partial_{\mu}f(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})\right],\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t\right.$
$\quad+\int_{0}^{T}\left\langle\nabla_{u}b(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})^{\top}\lambda_{t}^{1}+\nabla_{u}\sigma(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})^{\top}\lambda_{t}^{2},\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t$
$\quad+\int_{0}^{T}\left\langle\mathbb{E}^{\prime}\left[\partial_{\nu}b(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t},\widehat{\alpha}_{t})^{\top}{\lambda_{t}^{1}}^{\prime}+\partial_{\nu}\sigma(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})^{\top}{\lambda_{t}^{2}}^{\prime}\right],\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t$
$\quad-\left.\sum_{i\in I_{1}}\int_{[0,T]}\int_{0}^{t}\left\langle D_{\gamma}\phi^{i}(t,\widehat{X}_{t},\widetilde{\widehat{\alpha}}^{t})_{s},\alpha_{s}-\widehat{\alpha}_{s}\right\rangle\mathrm{d}s\mu^{i}(\mathrm{d}t)-\sum_{j\in I_{2}}\int_{0}^{T}\left\langle\nabla_{u}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta^{j}_{t},\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t\right]$
$\leq\mathbb{E}\left[\int_{0}^{T}r_{0}\left\langle\nabla_{u}f(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})+\mathbb{E}^{\prime}\left[\partial_{\mu}f(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})\right],\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t\right.$
$\quad+\int_{0}^{T}\left\langle\nabla_{u}b(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})^{\top}\lambda_{t}^{1}+\nabla_{u}\sigma(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})^{\top}\lambda_{t}^{2},\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t$
$\quad+\int_{0}^{T}\left\langle\mathbb{E}^{\prime}\left[\partial_{\nu}b(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t},\widehat{\alpha}_{t})^{\top}{\lambda_{t}^{1}}^{\prime}+\partial_{\nu}\sigma(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})^{\top}{\lambda_{t}^{2}}^{\prime}\right],\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t$ (42)
$\quad-\left.\sum_{i\in I_{1}}\int_{0}^{T}\left\langle\int_{[t,T]}D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}\mu^{i}(\mathrm{d}s),\alpha_{s}-\widehat{\alpha}_{s}\right\rangle\mathrm{d}t-\sum_{j\in I_{2}}\int_{0}^{T}\left\langle\nabla_{u}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta^{j}_{t},\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t\right].$
where we have used the Fubini theorem and the fact that the integral $\int_{t}^{T}D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}\mu^{i}(\mathrm{d}s)$ can be viewed as a Bochner integral valued in $\mathbb{R}^{l}$. Here, $D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}$ is the evaluation of the continuous function $D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})$ at time $t$. Then, the arbitrariness of $\alpha\in\mathscr{U}[0,T]$ in (3.3) yields that, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.
$\left\langle\mathbb{E}^{\prime}\left[r_{0}\partial_{\nu}f(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})+\partial_{\nu}b(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t},\widehat{\alpha}_{t})^{\top}{\lambda_{t}^{1}}^{\prime}+\partial_{\nu}\sigma(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t},\widehat{\alpha}_{t})^{\top}{\lambda_{t}^{2}}^{\prime}\right]\right.$
$\qquad+\left.r_{0}\nabla_{u}f(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})+\nabla_{u}b(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})^{\top}\lambda_{t}^{1}+\nabla_{u}\sigma(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})^{\top}\lambda_{t}^{2}\right.$
$\qquad-\left.\sum_{i\in I_{1}}\int_{[t,T]}D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}\mu^{i}(\mathrm{d}s)-\sum_{j\in I_{2}}\nabla_{u}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta_{t}^{j},u-\widehat{\alpha}_{t}\right\rangle\geq 0,\quad\forall u\in U.$ (43)
Using the representation (35) of the Hamiltonian $H^{r_{0}}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathbb{R}^{l}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\mapsto\mathbb{R}$, we obtain the stochastic minimum condition (3.11) from (3.3) by noting that $\lambda^{1}=Y$ and $\lambda^{2}=Z$ where the process pair $(Y,Z)=(Y_{t},Z_{t})_{t\in[0,T]}$ satisfies the BSDE (3.11).
On the other hand, the Radon measure $\mu^{i}$ is supported on $\{t\in[0,T];~{}\mathbb{E}[\phi^{i}(t,\widehat{X}_{t},\widetilde{\widehat{\alpha}}^{t})]=0\}$ for $i\in I_{1}$, which is a consequence of the first complementary slackness condition given in (33) of Lemma 3.10. Moreover, $\eta_{t}^{j}$ vanishes on the set $\{t\in[0,T];~{}\psi^{j}(t,\widehat{X}_{t}(\omega),\widehat{\alpha}_{t}(\omega))\neq 0\}$, $\mathrm{d}t$-$\mathrm{a.s.}$, which results from the second complementary slackness condition in (33) of Lemma 3.10. In addition, we have that $r_{0}+\sum_{i\in I_{1}}\mu^{i}([0,T])+\sum_{j\in I_{2}}\|\eta^{j}\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}=r_{0}+\sum_{i\in I_{1}}\|\mu^{i}\|_{\rm TV}+\sum_{j\in I_{2}}\|\eta^{j}\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}=1$ as $\mu^{i}\in\mathcal{M}^{+}[0,T]$ for $i\in I_{1}$. Finally, the uniqueness of solution to BSDE (3.11) is a result of the classical BSDE theory (c.f. Pham [37]). ∎
Remark 3.13.
According to the proof of Theorem 3.11, the adjoint process $(Y,Z)=(Y_{t},Z_{t})_{t\in[0,T]}$ in the constrained BSDE (3.11) can be actually interpreted as the components of the multiplier with respect to the McKean-Vlasov SDE (equality) constraint.
Remark 3.14.
In fact, Theorem 3.11 can be also applied to C-(E)MFC problem with Poisson jumps, which is formulated as follows:
$\begin{cases}\text{{\rm minimize} $J(\alpha)$ {\rm over} $\alpha\in\mathscr{U}[0,T];$}\\[6.00006pt] \displaystyle\text{{\rm s.t.} $X_{t}^{\alpha}=\kappa+\int_{0}^{t}b(s,X_{s}^{\alpha},\mathcal{L}(X_{s}^{\alpha},\alpha_{s}),\alpha_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,X_{s}^{\alpha},\mathcal{L}(X_{s}^{\alpha},\alpha_{s}),\alpha_{s})\mathrm{d}W_{s}$}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$+\int_{0}^{t}\int_{Z}\gamma(s,X_{s}^{\alpha},\mathcal{L}(X_{s}^{\alpha},\alpha_{s}),\alpha_{s},z)\tilde{N}(\mathrm{d}s,\mathrm{d}z)$,~{}~{}$\forall t\in[0,T]$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\mathbb{E}\left[\phi^{i}\left(t,X_{t}^{\alpha},\widetilde{\alpha}^{t}\right)\right]\geq 0$,~{}~{}$i\in I_{1}$,~{}~{}$\forall t\in[0,T]$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\psi^{j}\left(t,X_{t}^{\alpha},\alpha_{t}\right)\geq 0$,~{}~{}$j\in I_{2}$,~{}~{}$\mathrm{d}t\times\mathrm{d}\mathbb{P}$-{\rm a.s}.}\end{cases}$ (44)
Here, $\tilde{N}(\mathrm{d}s,\mathrm{d}z):=N(\mathrm{d}s,\mathrm{d}z)-\lambda_{Z}(\mathrm{d}z)\mathrm{d}s$ is a compensated Poisson random measure on some measurable space $(Z,\mathscr{Z})$ with finite intensity measure $\lambda_{Z}(\mathrm{d}z)\mathrm{d}s$, and the jump coefficient $\gamma:[0,T]\times\mathbb{R}^{n}\times{\cal P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathbb{R}^{l}\times Z\mapsto\mathbb{R}^{n}$ is assumed to satisfy
  • (B1)
    $\gamma$ is borel measurable and continuously differentiable w.r.t. $(x,u)\in\mathbb{R}^{n}\times\mathbb{R}^{l}$. Furthermore, $\gamma$ admits a continuous L-derivative w.r.t. the measure $\rho\in{\cal P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})$.
  • (B2)
    there is a constant $M>0$ such that, for all $(x,\rho),(x^{\prime},\rho^{\prime})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})$ and $(t,u,z)\in[0,T]\times\mathbb{R}^{l}\times Z$,
    $\left|\gamma(t,x^{\prime},\rho^{\prime},u,z)-\gamma(t,x,\rho,u,z)\right|$ $\leq M\left(|x-x^{\prime}|+\mathcal{W}_{2}(\rho,\rho^{\prime})\right),$
    $\left|\gamma(t,x,\rho,u,z)\right|+\left|\nabla_{u}\gamma(t,x,\rho,u,z)\right|$ $\leq M\left(1+|x|+|u|+M_{2}(\rho)\right).$
Correspondingly, the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ is defined as the natural filtration generated by the Brownian motion $W$ and the Poisson random measure $N(\mathrm{d}s,\mathrm{d}z)$. Then, the space $\mathbb{M}^{2}$ corresponding to the McKean-Vlasov jump diffusion state process in (44) should be modified to the set of $\mathbb{F}$-adapted (right continuous with left limits) $\mathbb{R}^{n}$-valued process $\eta=(\eta_{t})_{t\in[0,T]}$ admitting the following Itô’s representation:
$\eta_{t}=\eta_{0}+\int_{0}^{t}\eta^{1}_{s}\mathrm{d}s+\int_{0}^{t}\eta_{s}^{2}\mathrm{d}W_{s}+\int_{0}^{t}\int_{Z}\eta_{s}^{3}(z)\tilde{N}(\mathrm{d}s,\mathrm{d}z),$ (45)
where $\eta_{0}\in L^{2}(\Omega;\mathbb{R}^{n})$ is $\mathcal{F}_{0}$-measurable, $\eta^{1}=(\eta_{t}^{1})_{t\in[0,T]}$ and $\eta^{2}=(\eta_{t}^{2})_{t\in[0,T]}$ taking values respectively on $\mathbb{R}^{n}$ and $\mathbb{R}^{n\times r}$ are $\mathbb{F}$-adapted processes such that $\mathbb{E}[\int_{0}^{T}(|\eta_{t}^{1}|^{2}+|\eta_{t}^{2}|^{2})\mathrm{d}t]<\infty$. Furthermore, the $\mathbb{R}^{n}$-valued process $\eta^{3}=(\eta_{t}^{3}(z))_{(t,z)\in[0,T]\times Z}$ is $\mathbb{F}$-adapted and satisfies that $\mathbb{E}[\int_{0}^{T}\int_{Z}|\eta^{3}_{t}(z)|^{2}\lambda_{Z}(\mathrm{d}z)\mathrm{d}t]<\infty$. Thus, $\mathbb{M}^{2}$ becomes a Hilbert space under the modified inner product $\langle\cdot,\cdot\rangle_{\mathbb{M}^{2}}$ defined by, for all $\eta,\zeta\in\mathbb{M}^{2}$,
$\langle\eta,\zeta\rangle_{\mathbb{M}^{2}}$ $:=\mathbb{E}\left[\langle\eta_{0},\zeta_{0}\rangle+\int_{0}^{T}\left(\langle\eta_{t}^{1},\zeta_{t}^{1}\rangle+\mathrm{tr}\left(\eta_{t}^{2}(\zeta_{t}^{2})^{\top}\right)\right)\mathrm{d}t+\int_{0}^{T}\int_{Z}\langle\eta^{3}_{t}(z),\zeta^{3}_{t}(z)\rangle\lambda_{Z}(\mathrm{d}z)\mathrm{d}t\right].$ (46)
Note that Lemma 3.5 still holds because $\mathcal{L}(X_{t})$ is continuous (c.f. [9]). Moreover, Lemma 3.8 and Lemma 3.9 in this jump-diffusion framework are also valid. One can deduce the SMP for problem (44) by following the proof of Theorem 3.11: there exist some $r_{0}\geq 0$, $\mu^{i}\in\mathcal{M}^{+}[0,T]$ for $i\in I_{1}$ and $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $j\in I_{2}$ such that conditions (i)-(iii) of Theorem 3.11 hold and the stochastic (first-order) minimum condition is given by, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-$\mathrm{a.s.}$,
$\left\langle\nabla_{u}H^{r_{0}}(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t},p_{t},P_{t},K_{t})+\mathbb{E}^{\prime}\left[\partial_{\nu}H^{r_{0}}(t,\widehat{X}_{t}^{\prime},\widehat{\rho}_{t},\widehat{\alpha}_{t}^{\prime},p_{t},P_{t},K_{t})(\widehat{X}_{t},\widehat{\alpha}_{t})\right]\right.$
$\qquad\left.-\sum_{i\in I_{1}}\int_{t}^{T}D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}\mu^{i}(\mathrm{d}s)-\sum_{j\in I_{2}}\nabla_{u}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta_{t}^{j},u-\widehat{\alpha}_{t}\right\rangle\geq 0,\quad\forall u\in U.$ (47)
The Hamiltonian $H^{r_{0}}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathbb{R}^{l}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\times L^{2}((Z,\mathscr{Z},\lambda_{Z});\mathbb{R}^{n})\mapsto\mathbb{R}$ is defined by
$H^{r_{0}}(t,x,\rho,u,p,P,K)$ $:=\langle b(t,x,\rho,u),y\rangle+\mathrm{tr}\left(\sigma(t,x,\rho,u)z^{\top}\right)+\int_{Z}\gamma(t,x,\rho,u,z)K(z)\lambda_{Z}(\mathrm{d}z)$
$\quad+r_{0}f(t,x,\rho,u),$ (48)
and the process triplet $(p,P,K)=(p_{t},P_{t},K_{t})_{t\in[0,T]}$ taking values in $\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\times L^{2}((Z,\mathscr{Z},\lambda_{Z});\mathbb{R}^{n})$ is the unique solution to the following BSDE:
$p_{t}$ $=r_{0}\nabla_{x}g(\widehat{X}_{T},\widehat{\mu}_{T})+r_{0}\mathbb{E}^{\prime}\left[\partial_{\mu}g(\widehat{X}_{T},\widehat{\mu}_{T})(\widehat{X}_{T}^{\prime})\right]+\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\rho}_{s},\widehat{\alpha}_{s},p_{s},P_{s},K_{s})\mathrm{d}s$
$\quad+\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}H^{r_{0}}(s,\widehat{X}_{s}^{\prime},\widehat{\rho}_{s},\widehat{\alpha}_{s}^{\prime},p_{s},P_{s},K_{s})(\widehat{X}_{s},\widehat{\alpha}_{s})\right]\mathrm{d}s-\sum_{i\in I_{1}}\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)$
$\quad-\sum_{j\in I_{2}}\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\alpha}_{s})\mathrm{d}A_{s}^{j}-\int_{t}^{T}Z_{s}\mathrm{d}W_{s}-\int_{t}^{T}\int_{Z}K_{s}(z)\tilde{N}(\mathrm{d}s,\mathrm{d}z).$ (49)
In particular, when $m=0$, similar to Remark 3.12, the above BSDE (3.14) reduces to Eq. (36) in [9] in the McKean-Vlasov jump-diffusion model.
4 Constrained Stochastic Control and MFG Problems
4.1 Constrained stochastic control problem
In this subsection, we consider the following classical constrained stochastic control problem as a special case of MFC problem (22) given by
$\begin{cases}\text{minimize $J(\alpha)$ over $\alpha\in\mathscr{U}[0,T]$;}\\[6.00006pt] \displaystyle\text{subject to $X_{t}^{\alpha}=\kappa+\int_{0}^{t}b(s,X_{s}^{\alpha},\alpha_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,X_{s}^{\alpha},\alpha_{s})\mathrm{d}W_{s}$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\mathbb{E}\left[\phi^{i}\left(t,X_{t}^{\alpha},\widetilde{\alpha}^{t}\right)\right]\geq 0$,~{}~{}$i\in I_{1}$,~{}~{}$\forall t\in[0,T]$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\psi^{j}\left(t,X_{t}^{\alpha},\alpha_{t}\right)\geq 0$,~{}~{}$j\in I_{2}$,~{}~{}$\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s..}\end{cases}$ (50)
Then, we have from Theorem 3.11 that
Corollary 4.1 (SMP for constrained stochastic control problem).
Let $(\widehat{X},\widehat{\alpha})\in C:=\mathbb{M}^{2}\times\mathscr{U}[0,T]$ be an optimal solution to the constrained stochastic control problem (50). Then, there exist some $r_{0}\geq 0$, $\mu^{i}\in\mathcal{M}^{+}[0,T]$ for $i\in I_{1}$ and $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $j\in I_{2}$ such that
  • (i)
    the positive Radon measure $\mu^{i}$ is supported on $\{t\in[0,T];~{}\mathbb{E}[\phi^{i}(t,\widehat{X}_{t},\widetilde{\widehat{\alpha}}^{t})]=0\}$ for all $i\in I_{1}$,
  • (ii)
    for $j\in I_{2}$, as $\eta_{t}^{j}\geq 0$, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-$\mathrm{a.s.}$ and $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$, we can define the (continuous) increasing process $A^{j}=(A_{t}^{j})_{t\in[0,T]}$ by $A_{t}^{j}=\int_{0}^{t}\eta_{s}^{j}\mathrm{d}s$. Then, for $\mathbb{P}$-$\mathrm{a.s.}~{}\omega$, the process $A^{j}(\omega)$ only increases on the set $\{t\in[0,T];~{}\psi^{j}(t,\widehat{X}_{t}(\omega),\widehat{\alpha}_{t}(\omega))=0\}$, $\mathrm{d}t$-$\mathrm{a.s.}$,
  • (iii)
    it holds that $r_{0}+\sum_{i\in I_{1}}\mu^{i}([0,T])+\sum_{j\in I_{2}}\|\eta^{j}\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}=1$.
Furthermore, the stochastic (first-order) minimum condition holds in the sense that, for any $u\in U$ and $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-$\mathrm{a.s.}$,
$\left\langle\nabla_{u}H^{r_{0}}(t,\widehat{X}_{t},\widehat{\alpha}_{t},Y_{t},Z_{t})-\sum_{i\in I_{1}}\int_{t}^{T}D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}\mu^{i}(\mathrm{d}s)-\sum_{j\in I_{2}}\nabla_{u}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta_{t}^{j},u-\widehat{\alpha}_{t}\right\rangle\geq 0.$ (51)
Here, the Hamiltonian $H^{r_{0}}:[0,T]\times\mathbb{R}^{n}\times\mathbb{R}^{l}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\mapsto\mathbb{R}$ is defined by
$H^{r_{0}}(t,x,u,y,z):=\langle b(t,x,u),y\rangle+\mathrm{tr}\left(\sigma(t,x,u)z^{\top}\right)+r_{0}f(t,x,u),$ (52)
and the process pair $(Y,Z)=(Y_{t},Z_{t})_{t\in[0,T]}$ taking values in $\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ is the unique solution to the BSDE:
$Y_{t}$ $=r_{0}\nabla_{x}g(\widehat{X}_{T})+\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s-\sum_{i\in I_{1}}\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)$ (53)
$\quad-\sum_{j\in I_{2}}\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\alpha}_{s})\mathrm{d}A_{s}^{j}-\int_{t}^{T}Z_{s}\mathrm{d}W_{s}.$
Consider the stochastic control problem with the dynamic expectation constraint and the dynamic state constraint specified by, for $i\in I_{1}$ and $j\in I_{2}$,
$\mathbb{E}[\phi^{i}(t,X_{t}^{\alpha})]\geq 0,~{}\forall t\in[0,T],~{}\text{and}~{}\psi^{j}(t,X_{t}^{\alpha})\geq 0,~{}~{}\mathrm{d}t\times\mathrm{d}\mathbb{P}\text{-}\mathrm{a.s.}$ (54)
Note that the constraint functions $\phi^{i}$ and $\psi^{j}$ in (54) are independent of the control process. Then, the stochastic (first-order) minimum condition (51) is reduced to that, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-$\mathrm{a.s.}$,
$\left\langle\nabla_{u}H^{r_{0}}(t,\widehat{X}_{t},\widehat{\alpha}_{t},Y_{t},Z_{t}),u-\widehat{\alpha}_{t}\right\rangle\geq 0,\quad\forall u\in U.$ (55)
Here, the process pair $(Y,Z)=(Y_{t},Z_{t})_{t\in[0,T]}$ taking values in $\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ is the unique solution to the following BSDE:
$Y_{t}$ $=r_{0}\nabla_{x}g(\widehat{X}_{T})+\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s-\sum_{i\in I_{1}}\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s})\mu^{i}(\mathrm{d}s)$
$\quad-\sum_{j\in I_{2}}\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s})\mathrm{d}A_{s}^{j}-\int_{t}^{T}Z_{s}\mathrm{d}W_{s}.$ (56)
In particular, Hu et al. [26] considered the optimal control of SDEs with the dynamic expectation constraints described by the 1st one in (54) (i.e., $m=1$ and $\phi:=\phi^{1}$). In this case, the stochastic (first-order) minimum condition (55) holds true. Correspondingly, the BSDE (4.1) turns to be
$Y_{t}$ $=r_{0}\nabla_{x}g(\widehat{X}_{T})+\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s-\int_{[t,T]}\nabla_{x}\phi(s,\widehat{X}_{s})\mu^{i}(\mathrm{d}s)$
$\quad-\int_{t}^{T}Z_{s}\mathrm{d}W_{s},$ (57)
which coincides with the result given in Hu et al. [26].
We can also consider the special case when $m=1$, $m_{1}=0$ and $\psi(t,x)=\psi^{1}(t,x)=x-a_{t}$ for $(t,x)\in[0,T]\times\mathbb{R}$. Here, $t\mapsto a_{t}$ is a deterministic continuous real-valued mapping. Thus, the constrained stochastic control problem (50) reduces to
$\begin{cases}\text{minimize $J(\alpha)$ over $\alpha\in\mathscr{U}[0,T]$;}\\[6.00006pt] \displaystyle\text{subject to $X_{t}^{\alpha}=\kappa+\int_{0}^{t}b(s,X_{s}^{\alpha},\alpha_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,X_{s}^{\alpha},\alpha_{s})\mathrm{d}W_{s}$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$X_{t}^{\alpha}\geq a_{t}$,~{}~{}$\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.}\end{cases}$ (58)
As a result of Corollary 4.1, the following stochastic (first-order) minimum condition holds, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-$\mathrm{a.s.}$, $\left\langle\nabla_{u}H^{r_{0}}(t,\widehat{X}_{t},\widehat{\alpha}_{t},Y_{t},Z_{t}),u-\widehat{\alpha}_{t}\right\rangle\geq 0$ for all $u\in U$, where the process pair $(Y,Z)=(Y_{t},Z_{t})_{t\in[0,T]}$ satisfies the constrained BSDE:
$\mathrm{d}Y_{t}=-\nabla_{x}H^{r_{0}}(t,\widehat{X}_{t},\widehat{\alpha}_{t},Y_{t},Z_{t})\mathrm{d}t+\mathrm{d}A_{t}+Z_{t}\mathrm{d}W_{t},\quad Y_{T}=\nabla_{x}g(\widehat{X}_{T}),$ (59)
where $A_{t}=\int_{0}^{t}\eta_{s}\mathrm{d}s$ is the adapted increasing process with $\eta=(\eta_{t})_{t\in[0,T]}$ being the multiplier given in Corollary 4.1 such that
$\mathbb{E}\left[\int_{0}^{T}(\widehat{X}_{t}-a_{t})\mathrm{d}A_{t}\right]=0.$
This also implies that, $\mathbb{P}$-$\mathrm{a.s.}$,
$\int_{0}^{T}(\widehat{X}_{t}-a_{t})\mathrm{d}A_{t}=0,$ (60)
in view of $\widehat{X}_{t}-a_{t}\geq 0$, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.. Therefore, we can see that Eq. (59) resembles a reflected BSDE with the reflecting barrier $t\mapsto a_{t}$.
4.2 Constrained (extended) mean field game
In this section, we discuss the case of C-(E)MFG problem. We interpret the constrained MFG problem as a constrained stochastic control problem with the fixed point consistency condition.
We formulate the C-(E)MFG as follows:
$\begin{cases}\text{minimize $J(\alpha)$ over $\alpha\in\mathscr{U}[0,T]$;}\\[6.00006pt] \displaystyle\text{subject to $X_{t}^{\alpha}=\kappa+\int_{0}^{t}b(s,X_{s}^{\alpha},\rho_{s},\alpha_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,X_{s}^{\alpha},\rho_{s},\alpha_{s})\mathrm{d}W_{s}$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\mathbb{E}\left[\phi^{i}\left(t,X_{t}^{\alpha},\widetilde{\alpha}^{t}\right)\right]\geq 0$,~{}~{}$i\in I_{1}$,~{}~{}$\forall t\in[0,T]$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\psi^{j}\left(t,X_{t}^{\alpha},\alpha_{t}\right)\geq 0$,~{}~{}$j\in I_{2}$,~{}~{}$\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\rho_{t}=\mathcal{L}(\widehat{X}_{t},\widehat{\alpha}_{t})$ for $t\in[0,T]$,}\end{cases}$ (61)
where the measure flow $\rho=(\rho_{t})_{t\in[0,T]}$ takes values in $\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})$. Here, $\widehat{\alpha}=(\widehat{\alpha}_{t})_{t\in[0,T]}$ is an (admissible) optimal control and $\widehat{X}=(\widehat{X}_{t})_{t\in[0,T]}$ is the resulting state process. The measure flow $\rho=(\rho_{t})_{t\in[0,T]}$ satisfying $\rho_{t}=\mathcal{L}(\widehat{X}_{t},\widehat{\alpha}_{t})$ for $t\in[0,T]$ is called a mean field equilibrium (MFE) of the C-(E)MFG problem (61).
Corollary 4.2 (SMP for C-(E)MFG problem).
Let $\rho_{t}$ be an MFE to the C-(E)MFG problem (61). Then, there exist some $r_{0}\geq 0$, $\mu^{i}\in\mathcal{M}^{+}[0,T]$ for $i\in I_{1}$ and $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $j\in I_{2}$ such that
  • (i)
    the positive radon measure $\mu^{i}$ is supported on $\{t\in[0,T];~{}\mathbb{E}[\phi^{i}(t,\widehat{X}_{t},\widetilde{\widehat{\alpha}}^{t})]=0\}$ for all $i\in I_{1}$,
  • (ii)
    for $j\in I_{2}$, as $\eta_{t}^{j}\geq 0$, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-$\mathrm{a.s.}$ and $\eta^{j}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$, we can define the (continuous) increasing process $A^{j}=(A_{t}^{j})_{t\in[0,T]}$ by $A_{t}^{j}=\int_{0}^{t}\eta_{s}^{j}\mathrm{d}s$. Then, for $\mathbb{P}$-$\mathrm{a.s.}~{}\omega$, the process $A^{j}(\omega)$ only increases on the set $\{t\in[0,T];~{}\psi^{j}(t,\widehat{X}_{t}(\omega),\widehat{\alpha}_{t}(\omega))=0\}$, $\mathrm{d}t$-$\mathrm{a.s.}$,
  • (iii)
    it holds that $r_{0}+\sum_{i\in I_{1}}\mu^{i}([0,T])+\sum_{j\in I_{2}}\|\eta^{j}\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}=1$.
Furthermore, the stochastic (first-order) minimum condition holds in the sense that, for any $u\in U$ and $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-$\mathrm{a.s.}$,
$\left\langle\nabla_{u}H^{r_{0}}(t,\widehat{X}_{t},\widehat{\alpha}_{t},\rho_{t},Y_{t},Z_{t})-\sum_{i\in I_{1}}\int_{t}^{T}D_{\gamma}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})_{t}\mu^{i}(\mathrm{d}s)-\sum_{j\in I_{2}}\nabla_{u}\psi^{j}(t,\widehat{X}_{t},\widehat{\alpha}_{t})\eta_{t}^{j},u-\widehat{\alpha}_{t}\right\rangle\geq 0,$ (62)
where the Hamiltonian $H^{r_{0}}$ is given in (35), and the process pair $(Y,Z)=(Y_{t},Z_{t})_{t\in[0,T]}$, taking values in $\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$, is the unique solution to the BSDE:
$Y_{t}$ $=r_{0}\nabla_{x}g(\widehat{X}_{T})+\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\alpha}_{s},\rho_{s},Y_{s},Z_{s})\mathrm{d}s-\sum_{i\in I_{1}}\int_{[t,T]}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widetilde{\widehat{\alpha}}^{s})\mu^{i}(\mathrm{d}s)$ (63)
$\quad-\sum_{j\in I_{2}}\int_{t}^{T}\nabla_{x}\psi^{j}(s,\widehat{X}_{s},\widehat{\alpha}_{s})\mathrm{d}A_{s}^{j}-\int_{t}^{T}Z_{s}\mathrm{d}W_{s}.$
Finally, the consistency condition holds that $\rho_{t}=\mathcal{L}(\widehat{X}_{t},\widehat{\alpha}_{t})$.
4.3 An example: dynamic constrained LQ-MFC problem
In this subsection, we present an example of linear-quadratic MFC problem under a type of dynamic mixed state-control constraint in which the optimal control can be characterized.
Consider the following one-dimensional controlled Mckean-Vlasov SDE that, for $\alpha\in\mathscr{U}[0,T]$ with control space $U=\mathbb{R}$,
$\mathrm{d}X_{t}^{\alpha}$ =$=\{b_{1}X_{t}^{\alpha}+b_{2}\mathbb{E}[X_{t}^{\alpha}]+b_{3}\alpha_{t}+b_{4}\mathbb{E}[\alpha_{t}]\}\mathrm{d}t+\{\sigma_{1}X_{t}^{\alpha}+\sigma_{2}\mathbb{E}[X_{t}^{\alpha}]+\sigma_{3}\alpha_{t}+\sigma_{4}\mathbb{E}[\alpha_{t}]\}\mathrm{d}W_{t},$ (64)
where $X_{0}=\kappa\in L^{2}((\Omega,\mathcal{F}_{0},\mathbb{P});\mathbb{R})$ and $W=(W_{t})_{t\in[0,T]}$ is a standard one-dimensional Brownian motion on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ with $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ being the natural filtration generated by $W$. Here, model parameters $b_{i},\sigma_{i}\in\mathbb{R}$ for $i=1,2,3,4$. The aim of the agent is to minimize the following cost functional, for cost parameters $q,\ell\geq 0$ and $v>0$,
$J(\alpha)=\frac{1}{2}\mathbb{E}\left[\int_{0}^{T}\{q(X_{t}^{\alpha})^{2}+vu_{t}^{2}\}\mathrm{d}t+\ell(X_{T}^{\alpha}-\mathbb{E}[X_{T}^{\alpha}])^{2}\right],$ (65)
under dynamic state-control constraint described by, $\mathbb{P}$-a.s.
$h_{t}X_{t}^{\alpha}-\alpha_{t}\geq 0,\quad\forall t\in[0,T]$ (66)
with $t\mapsto h_{t}$ being a continuous function on $[0,T]$.
We next give a necessary optimality condition for constrained control problem (64)-(66) using the KKT condition that is established in Corollary 2.13. To this end, let $(\widehat{X},\widehat{\alpha})\in C=\mathbb{M}^{2}\times\mathscr{U}[0,T]$ be an optimal pair. First, we verify that LICQ(($\widehat{X},\widehat{\alpha}$)) holds true. In fact, by using Lemma 3.7, it is enough to show that the following mapping is surjective:
$\left(DF(\widehat{X},\widehat{\alpha}),D\Psi(\widehat{X},\widehat{\alpha})\right):\mathbb{M}^{2}\times\mathbb{L}_{\mathbb{F},\mathbb{R}}^{2}\to\mathbb{M}^{2}\times\mathbb{L}_{\mathbb{F},\mathbb{R}}^{2},$
where, for $(L,K)\in\mathbb{M}^{2}\times\mathbb{L}_{\mathbb{F},\mathbb{R}}$,
$D\Psi(\widehat{X},\widehat{\alpha})(L,K)$ $=-h(\cdot)L_{t}+K_{\cdot},$
$DF(\widehat{X},\widehat{\alpha})(L,K)$ $=L_{\cdot}-\int_{0}^{\cdot}b_{1}L_{s}\mathrm{d}s-\int_{0}^{\cdot}\sigma_{1}L_{s}\mathrm{d}W_{s}-\int_{0}^{\cdot}b_{2}\mathbb{E}[L_{s}]\mathrm{d}s-\int_{0}^{\cdot}\sigma_{2}\mathbb{E}[L_{s}]\mathrm{d}W_{s}$
$\quad-\int_{0}^{\cdot}b_{3}K_{s}\mathrm{d}s-\int_{0}^{\cdot}\sigma_{3}K_{s}\mathrm{d}W_{s}-\int_{0}^{\cdot}b_{4}\mathbb{E}[K_{s}]\mathrm{d}s-\int_{0}^{\cdot}\sigma_{4}\mathbb{E}[K_{s}]\mathrm{d}W_{s}.$
However, for any $(L^{\prime},K^{\prime})\in\mathbb{M}^{2}\times\mathbb{L}_{\mathbb{F},\mathbb{R}}^{2}$, we set $K_{t}=h_{t}L_{t}+K_{t}^{\prime}$ for $t\in[0,T]$, and plug it into the above equality. Then, there exists some $L\in\mathbb{M}^{2}$ guaranteed by Lemma 3.8 such that
$\left(D\Psi(\widehat{X},\widehat{\alpha}),DF(\widehat{X},\widehat{\alpha})\right)(L,K)=(L^{\prime},K^{\prime}).$
Thus, from Corollary 2.13 and Theorem 3.11, it follows that the KKT condition holds. Therefore, there exists some $\eta\in\mathbb{L}_{\mathbb{F},\mathbb{R}+}^{2}$ such that
$\eta_{t}\left(h_{t}\widehat{X}_{t}-\widehat{\alpha}_{t}\right)=0,\quad\mathrm{d}t\times\mathrm{d}\mathbb{P}\text{-}\mathrm{a.s.}.$ (67)
The constrained BSDE given in Theorem 3.11 is specified into
$\begin{cases}\displaystyle\mathrm{d}Y_{t}=-\left\{b_{1}Y_{t}+b_{2}\mathbb{E}[Y_{t}]+\sigma_{1}Z_{t}+\sigma_{2}\mathbb{E}[Z_{t}]+q\widehat{X}_{t}\right\}\mathrm{d}t+h_{t}\eta_{t}\mathrm{d}t+Z_{t}\mathrm{d}W_{t},\\[8.00003pt] \displaystyle Y_{T}=\ell\{\widehat{X}_{T}-\mathbb{E}[\widehat{X}_{T}]\}.\end{cases}$ (68)
Then, the stochastic minimal condition is given by, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.
$\{b_{3}Y_{t}+b_{4}\mathbb{E}[Y_{t}]+\sigma_{3}Z_{t}+\sigma_{4}\mathbb{E}[Z_{t}]+v\widehat{\alpha}_{t}+\eta_{t}\}(u-\widehat{\alpha}_{t})\geq 0,\quad\forall u\in U=\mathbb{R},$
which yields that the optimal control is characterized by
$\widehat{\alpha}_{t}=-v^{-1}\left\{b_{3}Y_{t}+b_{4}\mathbb{E}[Y_{t}]+\sigma_{3}Z_{t}+\sigma_{4}\mathbb{E}[Z_{t}]+\eta_{t}\right\},\quad\forall t\in[0,T].$ (69)
We also note that, if one considers problem (64)-(65) without constraint (66), the multiplier $\eta\in\mathbb{L}_{\mathbb{F},\mathbb{R}+}^{2}$ becomes zero. Then, the condition (67) trivially holds. Thus, the BSDE given in Theorem 3.11 becomes the classical one that
$\begin{cases}\displaystyle\mathrm{d}Y_{t}=-\left\{b_{1}Y_{t}+b_{2}\mathbb{E}[Y_{t}]+\sigma_{1}Z_{t}+\sigma_{2}\mathbb{E}[Z_{t}]+q\widehat{X}_{t}\right\}\mathrm{d}t+Z_{t}\mathrm{d}W_{t},\\[8.00003pt] \displaystyle Y_{T}=\ell\{\widehat{X}_{T}-\mathbb{E}[\widehat{X}_{T}]\}.\end{cases}$ (70)
The stochastic minimal condition is given by, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.
$\{b_{3}Y_{t}+b_{4}\mathbb{E}[Y_{t}]+\sigma_{3}Z_{t}+\sigma_{4}\mathbb{E}[Z_{t}]+v\widehat{\alpha}_{t}\}(u-\widehat{\alpha}_{t})\geq 0,\quad\forall u\in U=\mathbb{R}.$
This yields that the optimal control without constraint is
$\widehat{\alpha}_{t}=-v^{-1}\left\{b_{3}Y_{t}+b_{4}\mathbb{E}[Y_{t}]+\sigma_{3}Z_{t}+\sigma_{4}\mathbb{E}[Z_{t}]\right\},\quad\forall t\in[0,T],$ (71)
which coincides with the result for the extended MFC problem in the LQ framework as studied in [9] and an analytic solution can be derived by assuming that $Y_{t}$ takes the form $\beta_{t}X_{t}+\zeta_{t}\mathbb{E}[X_{t}]$ and solving the corresponding Riccati equation similar to [9].
Acknowledgements L. Bo and J. Wang are supported by Natural Science Basic Research Program of Shaanxi (No. 2023-JC-JQ-05), Shaanxi Fundamental Science Research Project for Mathematics and Physics (No. 23JSZ010) and Fundamental Research Funds for the Central Universities (No. 20199235177). X. Yu is supported by the Hong Kong RGC General Research Fund (GRF) under grant no. 15306523 and grant no. 15211524.
Appendix A Proofs of Auxiliary Results
In this appendix, we collect proofs of some auxiliary results in the main body of the paper.
Proof of Lemma 3.5.
Let $i\in I_{1}$. Then, for any $t_{1},t_{2}\in[0,T]$ with $t_{1}<t_{2}$, it holds that
$|g_{i}(t_{2})-g_{i}(t_{1})|$ $\leq\mathbb{E}[|\phi^{i}(t_{1},X_{t_{1}},\widetilde{\alpha}^{t_{1}})-\phi^{i}(t_{2},X_{t_{2}},\widetilde{\alpha}^{t_{2}})|]$
$\leq\mathbb{E}[|\phi^{i}(t_{1},X_{t_{1}},\widetilde{\alpha}^{t_{1}})-\phi^{i}(t_{2},X_{t_{2}},\widetilde{\alpha}^{t_{1}})|]+\mathbb{E}[|\phi^{i}(t_{2},X_{t_{2}},\widetilde{\alpha}^{t_{1}})-\phi^{i}(t_{2},X_{t_{2}},\widetilde{\alpha}^{t_{2}})|]$
$:=v_{1}+v_{2}.$
Note that $v_{1}$ converges to 00 as $|t_{1}-t_{2}|\to 0$ by DCT and the continuity of $\phi^{i}$ in $(t,x)$. Furthermore, by DCT, $\mathbb{E}[\|\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}\|_{\mathcal{X}}^{2}]=\mathbb{E}[\int_{t_{1}}^{t_{2}}|\alpha_{s}|^{2}\mathrm{d}s]$ converges to 00 as $|t_{1}-t_{2}|\to 0$ since $\alpha\in\mathbb{L}_{\mathbb{F}}^{2}$. For the term $v_{2}$, we have that, as $|t_{1}-t_{2}|\to 0$,
$v_{2}$ $=\mathbb{E}[|\langle D_{\gamma}\phi^{i}(t_{2},X_{t_{2}},\widetilde{\alpha}^{t_{1}}),\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}\rangle_{\mathcal{X}}|]+o(\mathbb{E}[\|\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}]\|_{\mathcal{X}})$
$\leq\left(\mathbb{E}[\|D_{\gamma}\phi^{i}(t_{2},X_{t_{2}},\widetilde{\alpha}^{t_{1}})\|_{\mathcal{X}}^{2}]\right)^{\frac{1}{2}}\left(\mathbb{E}[\|\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}\|_{\mathcal{X}}^{2}]\right)^{\frac{1}{2}}+o(\left(\mathbb{E}[\|\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}\|_{\mathcal{X}}^{2}]\right)^{\frac{1}{2}}).$
By using Assumption 1-(A2), $\mathbb{E}[\|D_{\gamma}\phi^{i}(t_{2},X_{t_{2}},\widetilde{\alpha}^{t_{1}})\|_{\mathcal{X}}^{2}]<\infty$, and hence $v_{2}\to 0$ as $|t_{1}-t_{2}|\to 0$ due to the fact $\mathbb{E}[\|\widetilde{\alpha}^{t_{2}}-\widetilde{\alpha}^{t_{1}}\|_{\mathcal{X}}^{2}]=\mathbb{E}[\int_{t_{1}}^{t_{2}}|\alpha_{s}|^{2}\mathrm{d}s]$. ∎
Proof of Lemma 3.8.
For any $Z\in\mathbb{L}_{\mathbb{F}}^{2}$, the classical SDE theory (c.f. Karatzas and Shreve [28]) gives that the following SDE has a unique strong solution $Y\in\mathbb{L}_{\mathbb{F}}^{2}$:
$\mathrm{d}Y_{t}=B(t,\omega,Y_{t},\overline{Z}_{t})\mathrm{d}t+\Sigma(t,\omega,Y_{t},\overline{Z}_{t}),\quad Y_{0}\in L^{2}((\Omega,\mathcal{F}_{0},\mathbb{P});\mathbb{R}^{n}),$
where, for $t\in[0,T]$, $\overline{Z}_{t}$ stands for the r.v. $Z_{t}$ instead of the evaluation $Z_{t}(\omega)$ at $\omega\in\Omega$. Then, we have constructed a self-mapping $G:\mathbb{L}_{\mathbb{F}}^{2}\mapsto\mathbb{L}_{\mathbb{F}}^{2}$ by $G(Z)=Y$. Thus, the existence of solution reduces to showing that the mapping $G$ admits a fixed point. For $i=1,2$, let $Y^{i}=(Y_{t}^{i})_{t\in[0,T]}$ be the corresponding solution associated to $Z^{i}\in\mathbb{L}_{\mathbb{F}}^{2}$. Let $M=M(T)>0$ be a generic constant that may differ from line to line. The standard moment estimation under the assumption (28) on coefficients $(B,\Sigma)$ results in
$\mathbb{E}\left[|Y_{t}^{1}-Y_{t}^{2}|^{2}\right]\leq M\int_{0}^{t}\left\{\mathbb{E}\left[|Y_{s}^{1}-Y_{s}^{2}|^{2}\right]+\mathbb{E}\left[|Z_{s}^{1}-Z_{s}^{2}|^{2}\right]\right\}\mathrm{d}s,~{}~{}\forall t\in[0,T].$
Applying the Gronwall’s inequality, we obtain that
$\mathbb{E}\left[|Y_{t}^{1}-Y_{t}^{2}|^{2}\right]\leq M\int_{0}^{t}\mathbb{E}\left[|Z_{s}^{1}-Z_{s}^{2}|^{2}\right]\mathrm{d}s,~{}~{}\forall t\in[0,T],$ (72)
which yields the continuity of $G:\mathbb{L}_{\mathbb{F}}^{2}\mapsto\mathbb{L}_{\mathbb{F}}^{2}$.
Next, let $Y^{0}_{t}=Y_{0}$ for all $t\in[0,T]$, and define $Y_{t}^{n}$ $(n\in\mathbb{Z}_{+})$ by induction as follows:
$Y_{t}^{n}=Y_{0}+\int_{0}^{T}B\left(t,\omega,Y_{t}^{n},\overline{Y_{t}^{n-1}}\right)\mathrm{d}t+\int_{0}^{T}\Sigma\left(t,\omega,Y_{t}^{n},\overline{Y_{t}^{n-1}}\right)\mathrm{d}W_{t}.$ (73)
Under the assumption (28), it holds from (72) that, for all $n\in\mathbb{N}$,
$\mathbb{E}\left[|Y_{t}^{n+1}-Y_{t}^{n}|^{2}\right]\leq M\int_{0}^{t}\mathbb{E}\left[|Y_{t}^{n}-Y_{t}^{n-1}|^{2}\right]\mathrm{d}s,~{}~{}\forall t\in[0,T].$
This implies that $|Y^{n+1}-Y^{n}|_{\mathbb{L}_{\mathbb{F}}^{2}}^{2}=O(\frac{1}{n!})$ as $n\to\infty$. Thus, $(Y_{n})_{n\in\mathbb{N}}$ is a Cauchy sequence in $\mathbb{L}_{\mathbb{F}}^{2}$, and hence there exists a $Y\in\mathbb{L}_{\mathbb{F}}^{2}$ such that $Y^{n}\to Y$, as $n\to\infty$, in $\mathbb{L}_{\mathbb{F}}^{2}$. Letting $n\to\infty$, we can see that $Y\in\mathbb{L}_{\mathbb{F}}^{2}$ is a fixed point of the mapping $G:\mathbb{L}_{\mathbb{F}}^{2}\mapsto\mathbb{L}_{\mathbb{F}}^{2}$. By using (72) and the Gronwall’s inequality again, we conclude the uniqueness of the solution to the SDE (30). ∎